### v1

# How Real Are Synthetic Therapy Conversations? Evaluating Fidelity in Prolonged Exposure Dialogues

Abstract
The growing adoption of synthetic data in healthcare is driven by privacy concerns, limited access to real-world data, and the high cost of annotation. This work explores the use of synthetic Prolonged Exposure (PE) therapeutic conversations for Post-Traumatic Stress Disorder (PTSD) as a scalable alternative for training and evaluating clinical models. We systematically compare real and synthetic dialogues using linguistic, structural, and protocol-specific metrics, including turn-taking patterns and treatment fidelity. We also introduce and evaluate PE-specific metrics derived from linguistic analysis and semantic modeling, offering a novel framework for assessing clinical fidelity beyond surface fluency. Our findings show that although synthetic data holds promise for mitigating data scarcity and protecting patient privacy, it can struggle to capture the subtle dynamics of therapeutic interactions. In our dataset, synthetic dialogues match structural features of real-world dialogues (e.g., speaker switch ratio: 0.98 vs. 0.99), however synthetic interactions do not adequately reflect key fidelity markers (e.g., distress monitoring). We highlight gaps in existing evaluation frameworks and advocate for fidelity-aware metrics that go beyond surface fluency to uncover clinically significant failures. Our findings clarify where synthetic data can effectively complement real-world datasets—and where critical limitations remain.
How Real Are Synthetic Therapy Conversations?
Evaluating Fidelity in Prolonged Exposure Dialogues
Suhas BN1  Dominik Mattioli1  Saeed Abdullah1 Rosa I. Arriaga2  Chris W. Wiese3  Andrew M. Sherrill4 1College of Information Sciences and Technology, Penn State University, USA 2School of Interactive Computing, Georgia Tech, USA 3School of Psychology, Georgia Tech, USA 4Department of Psychiatry and Behavioral Sciences, Emory University, USA {bnsuhas,saeed}@psu.edu, andrew.m.sherrill@emory.edu
1 Introduction
Training machine learning models in sensitive domains like healthcare remains a challenge (Giuffrè and Shung, 2023; Kokosi and Harron, 2022). Real clinical conversations, particularly therapist-client dialogues, are rich sources of data for training models in areas such as mental health diagnostics, therapeutic interventions, and adherence monitoring. However, the high cost of data annotation, privacy concerns surrounding personally identifiable information (PII) (Kokosi and Harron, 2022; BN and Abdullah, 2022; BN et al., 2023), and ethical constraints in sharing sensitive interactions severely limit access to these datasets. Synthetic data has emerged as a promising alternative to address these challenges (Aher et al., 2023). By simulating conversations in a controlled environment, synthetic data offers scalability and privacy preservation while reducing the reliance on costly real-world annotations.
In trauma-focused mental health care, especially Prolonged Exposure (PE) therapy for PTSD, large language models offer the potential to generate synthetic therapy dialogues at scale. Despite its potential, questions remain about the fidelity and utility of synthetic data to capture the subtle dynamics of therapeutic interactions (e.g., emotional depth, turn-taking patterns, and building therapeutic alliances) (Shen et al., 2024). Recent work (e.g., BN et al. (2025)) has shown that synthetic dialogues can fluently simulate PE sessions across diverse contexts. However, such outputs may only mimic surface structure and tone. Subtle fidelity lapses—like premature therapist reflection or avoidance reinforcement—often remain undetected by automatic metrics or non-clinical annotators (Chiu et al., 2024; Zhang et al., 2024; Lee et al., 2024b).
For instance, a therapist saying “That’s a really powerful insight” mid-exposure may appear empathic but violates PE protocol by derailing trauma processing (see Fig. 1). Without clinical expertise, both humans and automated metrics tend to overestimate fidelity. This paper introduces a fidelity-focused lens to evaluate whether synthetic PE sessions truly adhere to the therapeutic process, not just structure or tone. This fidelity-focused lens involves evaluating dialogues not only on linguistic coherence but also on adherence to specific therapeutic protocols and the successful navigation of key clinical interaction patterns (e.g., managing avoidance, appropriate SUDS monitoring) using both automated metrics and expert clinical assessment.
In this study, we present a systematic comparison of real and synthetic clinical conversations, focusing on their linguistic, structural, and statistical characteristics. Our analysis explores whether synthetic data can replicate key features of real dialogues, including turn-taking patterns, protocol adherence, and emotional progression, while mitigating the risks associated with privacy and data scarcity. By analyzing both the strengths and limitations of synthetic therapy dialogues, our findings inform future improvements in the generation and evaluation of synthetic data for clinical applications. This study contributes to ongoing efforts to bridge the gap between real and synthetic data, providing a roadmap for the future of synthetic dialogue in sensitive domains. To our knowledge, this is the first study to combine statistical, linguistic, and clinical fidelity evaluations on both real and synthetic Prolonged Exposure (PE) therapy conversations at this scale.
2 Motivation
The development of synthetic clinical dialogues is motivated by several critical challenges in mental healthcare. Prolonged Exposure (PE) therapy—an evidence-based treatment for post-traumatic stress disorder (PTSD)—offers a structured therapeutic approach that could benefit from AI support, but faces significant obstacles:
  1. 1.
    Data scarcity: The limited availability of annotated clinical conversations restricts the development of AI systems for therapist training. Manual annotation is costly and time-consuming, often requiring up to a week per session per annotator, further compounding the scarcity of usable data.
  2. 2.
    Privacy constraints: Patient narratives contain highly sensitive personal information, making clinical data difficult to share or publish under privacy regulations and ethical guidelines.
  3. 3.
    Lack of diversity: Existing datasets often fail to capture diverse demographic profiles, trauma types, and therapeutic approaches, limiting model generalization.
  4. 4.
    Standardization challenges: In the absence of standardized benchmarks, fidelity evaluation of synthetic dialogues remains subjective and inconsistent.
Our work addresses these challenges by creating and systematically evaluating synthetic PE therapy conversations that maintain clinical fidelity while protecting patient privacy. Beyond privacy preservation, synthetic data enables scalable annotation, broader demographic and trauma-type coverage, and controlled variation in dialogue structure—capabilities difficult to achieve with curated real-world data alone. Our framework compares real and synthetic dialogues across linguistic, structural, and clinical fidelity dimensions.
3 Related Work
The evaluation of synthetic clinical conversation datasets, particularly for psychotherapy like Cognitive Behavioral Therapy (CBT) and general counseling, has grown significantly. However, research focusing on Prolonged Exposure (PE) therapy and other structured trauma-focused treatments remains limited. Existing studies primarily address (1) synthetic dataset generation and fidelity assessment, (2) evaluation metrics beyond lexical similarity, and (3) human-in-the-loop validation, often lacking PE-specific considerations.
3.1 Synthetic Dataset Generation and Evaluation
Several studies explore generating and assessing synthetic clinical datasets. For instance, Chiu et al. (2024) introduced BOLT to evaluate LLM-generated therapist behaviors (reflections, psychoeducation) in general counseling, while Qiu and Lan (2024) proposed SimPsyDial, benchmarking it against RealPsyDial using the Working Alliance Inventory (WAI-O-S). Zhang et al. (Zhang et al., 2024) presented CPsyCoun, reconstructing dialogues from reports and evaluating them with BERTScore (Zhang et al., 2019), GPTScore, and qualitative review. Others have focused on tailoring dialogues to specific counselor styles (Xie et al., 2024) or enhancing realism via data augmentation (Kim et al., 2024). Although datasets like Thousand Voices simulate structural variation (BN et al., 2025), current evaluations primarily compare synthetic data to real-world general counseling or CBT (Lee et al., 2024a; Shen et al., 2024). They generally lack explicit integration and assessment of trauma-specific markers or the structural fidelity crucial for PE, such as handling avoidance behaviors and managing imaginal exposure sequences. Our work extends these efforts by specifically evaluating whether synthetic PE sessions maintain clinical realism according to PE protocols (e.g., therapist containment, SUDS handling, avoidance redirection), highlighting fidelity violations where linguistic fluency masks procedural errors often missed by standard LLM evaluations or non-clinical reviewers.
3.2 Beyond Lexical Metrics
Evaluation is shifting from traditional lexical metrics (e.g., BLEU (Papineni et al., 2002)) towards more context-aware approaches. Transformer-based scores like BERTScore and GPTScore assess semantic alignment (Zhang et al., 2024; Xie et al., 2024), and psychological metrics (e.g., CTRS, PANAS in COUNSELINGEVAL (Lee et al., 2024a)) evaluate aspects like empathy and coherence. Specific interaction patterns, like Proactive Questioning Ability (PQA), have also been analyzed (Shen et al., 2024). However, these metrics are often unadapted for trauma-focused contexts, and advanced sequential techniques like Dynamic Time Warping (DTW), potentially useful for assessing turn-by-turn coherence in PE’s structured flow, remain largely unexplored in this domain. While qualitative assessments consider emotional tone and goal alignment (Zhang et al., 2024; Xie et al., 2024), and methods for structured questioning exist (Ren et al., 2024), robust quantitative frameworks for evaluating adherence to PE’s specific therapeutic structure are underdeveloped.
3.3 Human-in-the-Loop Validation and Engagement Metrics
Human expert validation is recognized as critical. Approaches like COUNSELINGEVAL (Lee et al., 2024a), PsyDT (Xie et al., 2024), and CPsyCoun (Zhang et al., 2024) incorporate clinician or professional reviewer feedback to assess quality, empathy, and safety. Concurrently, personalization and engagement metrics are emphasized (Lee et al., 2024a; Zhang et al., 2024; Xie et al., 2024), and prior work addresses empathy (Morris et al., 2018) and structured workflows (Ren et al., 2024). However, these human evaluations predominantly focus on general counseling or CBT contexts. They typically omit systematic assessment of PE-specific elements like trauma cue processing, therapeutic handling of emotional regulation, and adherence to PE protocols, especially when compared against real-world PE therapy data.
3.4 Bridging the Gap in PE Therapy Evaluation
Prolonged Exposure (PE) therapy is a first-line, empirically supported treatment for PTSD (Sherrill and Rauch, 2019). Despite progress in synthetic dialogue generation (Qiu and Lan, 2024; Zhang et al., 2024; Xie et al., 2024), evaluation metrics (Lee et al., 2024a; Zhang et al., 2024; Xie et al., 2024), and human validation (Lee et al., 2024a; Zhang et al., 2024; Xie et al., 2024), a significant gap persists in assessing the fidelity of synthetic data for structured, trauma-focused therapies like PE (Chiu et al., 2024; Qiu and Lan, 2024; Lee et al., 2024a; Shen et al., 2024; Zhang et al., 2024; Xie et al., 2024; Hu et al., 2024). As discussed, current approaches generally lack the specific focus and metrics needed to evaluate PE’s unique conversational structures (e.g., avoidance handling, trauma cue processing, imaginal exposure flow) and robustness to therapy-specific challenges like patient emotional dysregulation or engagement difficulties (Qiu and Lan, 2024; Lee et al., 2024a; Shen et al., 2024; Zhang et al., 2024; Xie et al., 2024; Kim et al., 2024). Our work directly addresses this limitation by systematically evaluating synthetic data alignment with structured PE interactions, identifying key areas for improvement in generating clinically valid PE therapy dialogues (See Table 1 and Appendix A).
Metric Yes No N/A
Therapist explained rationale for imaginal? $\square$ $\square$ $\square$
Therapist gave client instructions to carry out imaginal? $\square$ $\square$ $\square$
Hotspots procedure and rationale introduced? $\square$ $\square$ $\square$
Therapist helped patient to identify hotspots? $\square$ $\square$ $\square$
Therapist oriented the client to imaginal planned for that session? $\square$ $\square$ $\square$
Therapist monitored SUDS ratings about every 5 minutes? $\square$ $\square$ $\square$
Therapist used appropriate reinforcing comments during imaginal? $\square$ $\square$ $\square$
Therapist elicited thoughts and feelings as appropriate? $\square$ $\square$ $\square$
Therapist prompted for present tense, closed eyes? $\square$ $\square$ $\square$
Imaginal lasted about 30-45 minutes (or about 15 for final imaginal)? $\square$ $\square$ $\square$
Therapist processed the imaginal with client? $\square$ $\square$ $\square$
Table 1: Clinically validated fidelity metrics for therapist adherence to essential elements of imaginal exposure therapy (Foa et al., 2007; Powers et al., 2010; Rauch et al., 2009; Hembree et al., 2003). This checklist evaluates whether therapists consistently implement key procedural components, including providing rationale, guiding the client through the exposure process, monitoring distress levels, and reinforcing engagement. Each element is rated as ‘Yes,’ ‘No,’ or ‘N/A’ to ensure treatment fidelity, maintain therapeutic consistency, and identify areas for improvement in clinical practice.
4 Methodology
We analyzed 400 prolonged exposure therapy conversations—200 real-world and 200 synthetic. The synthetic data comes from the publicly available Thousand Voices of Trauma dataset (BN et al., 2025), generated using Claude Sonnet 3.5 (Anthropic, 2024) with PE-specific prompting frameworks adapted from clinical guidelines. This data includes diverse therapist-client interactions across trauma types, therapy phases, and demographics.
The real-world PE sessions were collected under IRB-approved protocols, with participant consent, and cannot be released due to privacy constraints. Each session (1–1.5 hours) was transcribed using Amazon HealthScribe (Services, 2023), manually verified, and reformatted by merging consecutive speaker turns for readability. Each conversation followed a validated therapy fidelity checklist (See Table 1) to align with real-world standards. Both datasets underwent the same preprocessing:
  1. 1.
    Standardized formatting for consistency.
  2. 2.
    Processing through ModernBERT (Warner et al., 2024) for analysis.
  3. 3.
    Removal of non-verbal cues (e.g., pauses, laughter) to focus on dialogue, with plans to incorporate these in future work on emotional speech.
All evaluation code, pre-processing pipelines, and metric definitions will be released to support reproducibility and external validation.
4.1 Metrics and Analysis
To systematically compare real and synthetic therapeutic conversations, we selected a diverse set of linguistic, structural, and statistical metrics. These metrics provide insights into conversational dynamics, protocol adherence, and overall fidelity, ensuring a holistic evaluation of synthetic dialogue generation. To reduce subjectivity, we adopt a fidelity checklist (Table 1) and PE-specific metrics (Sec. 5) grounded in existing clinical guidelines for evaluating therapeutic adherence. Our methodology consists of four key analyses: system-level metrics comparison, correlation analysis, statistical significance testing, and PE-specific metrics.
4.1.1 System-Level Metrics Comparison
We begin by measuring fundamental characteristics of the conversations, including turn-taking patterns, verbosity, lexical diversity, and readability (see Table 2). Key metrics include:
  1. 1.
    Turn-taking dynamics: Metrics such as Normalized Speaker Switches, Therapist-Client Turn Ratio, and Normalized Turn Duration capture the natural flow of conversation. These are essential for evaluating whether synthetic dialogues mimic real-world engagement.
  2. 2.
    Linguistic complexity and coherence: Average Utterance Length, Utterance Length Std Dev, and Readability Score assess how natural and readable the synthetic text is. A significant deviation from real data would indicate a lack of coherence in generated responses.
  3. 3.
    Lexical richness: Vocabulary Richness would help quantify lexical variety, providing insight into whether synthetic dialogues are overly repetitive.
  4. 4.
    Predictability of text: Flow Entropy and Perplexity can help measure randomness and fluency, determining whether the synthetic text is overly structured or unnatural.
    By comparing means and standard deviations, we evaluate whether synthetic conversations approximate real dialogue structures. Significant deviations signal potential areas where synthetic data falls short.
4.1.2 Correlation Analysis
While mean comparisons provide a general overview, correlation analysis (see Table 3) quantifies the consistency of relationships across metrics between real and synthetic data. This step helps us determine whether patterns observed in real data persist in synthetic dialogues. For instance:
  1. 1.
    Turn-taking consistency: Strong correlations in Normalized Therapist/Client Turns and Turn Ratios indicate that synthetic dialogues replicate real-world conversational structures.
  2. 2.
    Linguistic coherence: A high correlation in Utterance Length and Flow Entropy suggests that synthetic responses maintain natural variation in sentence length and randomness. Lower correlation values suggest that synthetic data may lack fidelity in preserving conversational nuances, particularly in aspects such as vocabulary richness and readability.
4.1.3 Statistical Significance Testing
To confirm whether differences between real and synthetic data are statistically meaningful, we conduct Mann-Whitney U tests (Table 4). This analysis assesses whether observed variations are due to randomness or inherent inconsistencies in the synthetic data. Key takeaways include:
  1. 1.
    Significant differences in linguistic complexity: Metrics such as Utterance Length, Turn Duration, and Vocabulary Richness show statistically significant differences, indicating that synthetic dialogues do not yet fully capture the expressive range of real conversations.
  2. 2.
    Minimal differences in turn-taking ratios: Therapist-Client Turn Ratio and Speaker Switches show weaker statistical significance, suggesting that structural patterns are relatively well-preserved.
  3. 3.
    Higher variability in generated text: The difference in Perplexity and Flow Entropy highlights challenges in maintaining linguistic variability without excessive repetition or over-simplification.
Figure 1: Selected annotated examples from synthetic PE therapy sessions. Each dialogue segment is evaluated for fidelity adherence based on PE protocol guidelines. Despite structural fluency, subtle violations like role drift (Ex. 1) and premature reflection (Ex. 4) highlight limitations in fidelity that escape automated scoring.
4.1.4 PE-Specific Metrics
Prolonged Exposure (PE) therapy reduces pathological fear through repeated trauma exposure. To assess the fidelity of synthetic PE sessions, we evaluate key therapeutic constructs (See Appendix A for definitions and Table 5): Trauma Narrative Coherence measures how structured and detailed a client’s trauma account is, reflecting cognitive integration. Emotional Engagement captures the level of emotional expression, linked to better outcomes. Avoidance Handling evaluates how well avoidance behaviors are addressed. Exposure Guidance assesses the therapist’s role in structuring effective exposure exercises. Cognitive Restructuring tracks how clients challenge maladaptive beliefs. Emotional Habituation and SUDS Progression measure distress reduction over repeated exposures. Avoidance Reduction quantifies improvements in engaging with trauma-related content. Emotion Intensity assesses the variability and magnitude of emotional responses. These metrics are derived using linguistic analysis, semantic modeling, and interaction patterns.
5 Findings
5.1 System-Level Metrics & Correlation
The system-level metrics comparison highlights alignment between real and synthetic dialogues, revealing both structural matches and model-driven differences. While turn-taking patterns (e.g., Normalized Speaker Switches, Therapist-Client Turn Ratio, and Normalized Therapist/Client Turns) remain similar, variations in utterance length, lexical diversity, and entropy-based measures arise due to prolonged exposure therapy and model constraints.
  1. 1.
    Turn-taking fidelity: Synthetic data closely aligns with real data in Normalized Speaker Switches (0.98 vs. 0.99) and Therapist-Client Turn Ratio (0.01 vs. 0.01), preserving conversational structure. Some deviations occur as real therapy sessions involve extended client turns, which LLMs struggle to maintain due to context limitations.
  2. 2.
    Concise and structured responses: Synthetic dialogues are shorter (Average Utterance Length: 22.90 ± 1.74 vs. 68.72 ± 26.61) and more consistent (Utterance Length Std Dev: 18.54 ± 2.35 vs. 135.85 ± 66.25) due to LLM output constraints. Techniques like Chain-of-Thought prompting helps improve coherence, though larger output contexts are needed.
  3. 3.
    Vocabulary Richness is marginally higher in synthetic data but may reflect repeated paraphrasing rather than authentic diversity. Real therapy involves longer client responses, naturally increasing lexical variety, but synthetic responses remain contextually appropriate.
  4. 4.
    Increased structural consistency: Higher Perplexity (21.22 vs. 14.73) and lower Flow Entropy (1.06 vs. 1.30) suggest a structured, and predictable flow. While real data exhibit spontaneity, synthetic dialogues maintain stability, benefiting structured evaluations.
  5. 5.
    Correlation analysis: High correlations in turn-taking metrics (0.65–0.78 synthetic vs. 0.85–0.92 real) confirm accurate conversational dynamics. Lower correlations in linguistic complexity metrics (e.g., Readability Score, Vocabulary Richness, Perplexity) reflect structural differences but do not hinder the coherence/naturalness of the conversation.
5.2 Statistical Significance Testing
The Mann-Whitney U test results confirm that many observed differences between real and synthetic dialogues are statistically significant ($p<0.05$). However, these differences stem from model design choices and practical constraints rather than fundamental shortcomings. Key observations include:
  1. 1.
    Distinct patterns in utterance structure: The test shows differences in utterance length, turn duration, and their standard deviations ($p<10^{-17}$). This is expected, as synthetic dialogues are designed to maintain coherence by producing more structured and concise responses. In real PE therapy, clients occasionally have extended monologues, which LLMs struggle to handle due to context window limitations. To compensate, we shorten utterances while preserving the conversational structure. Methods like Chain-of-Thought prompting have improved this, but achieving full parity would require larger output contexts.
  2. 2.
    Lexical properties follow a structured pattern: Differences in Vocabulary Richness ($p<10^{-15}$) and Flow Entropy ($p<10^{-17}$) indicate that synthetic data exhibits a more varied vocabulary but within a constrained framework. This is a direct result of the model prioritizing coherence and avoiding redundant expressions. While real conversations naturally contain more spontaneity, the synthetic approach ensures stability in generated dialogue while maintaining conversational depth.
  3. 3.
    Certain aspects remain comparable: Metrics such as Normalized Total Turns ($p=1.00$) and Readability Score ($p=0.28$) show no significant differences, meaning that despite shorter utterances, the number of conversational exchanges and overall readability remain aligned with real data. This suggests that while individual responses may be more concise, the overall flow and engagement in the conversation are well-preserved.
These findings reinforce that the synthetic model captures key conversational characteristics while ensuring structured, coherent responses. While differences exist, they align with known model constraints and do not compromise the overall integrity of the generated dialogues.
5.3 Feature Importance
Feature importance analysis highlights which metrics most influence synthetic data generation, reinforcing trends observed in prior evaluations. While not the primary focus, these insights provide valuable confirmation of the model’s alignment with conversational structure and linguistic patterns.
  1. 1.
    Turn-taking features contribute heavily: Metrics such as Normalized Speaker Switches and Therapist-Client Turn Ratio emerge as highly influential. This is a positive indication that the model effectively captures conversational flow, ensuring balanced exchanges between therapist and client. The strong correlation between real and synthetic data in these aspects suggests that the generated dialogues maintain natural interaction patterns.
  2. 2.
    Utterance length and entropy-based measures shape responses: Metrics like Average Utterance Length, Perplexity, and Flow Entropy stand out as key distinguishing factors. This aligns with statistical significance results, reflecting that synthetic responses are more structured and predictable. This behavior is expected, as the model prioritizes coherence and consistency, which helps maintain clarity even in complex dialogues. While synthetic responses may be more concise, they still reflect key interaction patterns but require further refinement to support emotionally dynamic exchanges.
  3. 3.
    Lexical richness plays a secondary role: Vocabulary Richness was moderately influential, indicating that synthetic data maintains a varied vocabulary while ensuring structured dialogue. While real-world conversations may exhibit greater linguistic flexibility, synthetic dialogues remain clear and contextually appropriate, balancing variation with coherence.
5.4 Prolonged Exposure Specific Metrics
Table 5 presents the statistical test results comparing synthetic Prolonged Exposure (PE) data with real therapy session data. The results indicate that the synthetic data effectively captures key therapeutic constructs, with several core PE metrics such as Trauma Narrative Coherence, Emotional Engagement, Avoidance Handling, Exposure Guidance, Cognitive Restructuring, Avoidance Reduction, and Emotion Intensity showing strong alignment with real data (p < 0.001) showing strong alignment with real data. These results suggest that the synthetic data successfully replicate the structure and engagement patterns seen in real therapy sessions. However, certain metrics do not reach statistical significance, indicating areas where synthetic data generation may need refinement. Emotional Habituation ($p=0.102$) and SUDS Progression ($p=0.073$) track distress reduction over repeated exposures, and their lack of significance suggests the synthetic data may not fully model this progressive decrease. Similarly, while Trauma Narrative Coherence is well-captured, Narrative Development ($p=0.251$) appears to be less robust, possibly due to limited context retention or rigid structure in generated responses. Despite these discrepancies, the overall performance of synthetic data suggests it is a viable alternative for privacy-sensitive applications. Addressing dynamic narrative evolution could further improve alignment with real data. However, fidelity violations in synthetic sessions are often subtle and may not disrupt structural metrics. For example, therapist utterances can appear empathetic or affirming while inadvertently shifting the session away from trauma anchoring. These moments often go unnoticed by automated scoring or non-clinical reviewers, underscoring the need for fidelity-aware evaluation frameworks that integrate human clinical judgment.
6 Discussion
In this paper, we evaluate the fidelity of synthetic Prolonged Exposure (PE) therapy conversations by comparing them to real interactions using linguistic, structural, and statistical analyses. We discuss the findings from four key perspectives: system-level metrics and correlation (Table 2), statistical significance testing (Table 3), feature importance (Table 4), and PE-specific metrics (Table 5).
6.1 Clinical Implications
Our findings have several important implications for clinical practice. First, synthetic PE sessions could serve as valuable training tools for novice therapists learning to identify protocol deviations, as they provide controlled examples of both adherent and non-adherent interactions without privacy concerns. Second, the consistent replication of structural elements (turn-taking, session flow) suggests synthetic data can supplement clinical training materials, particularly in regions with limited access to specialized trauma training. However, the identified limitations in capturing emotional habituation patterns and dynamic narrative development indicate synthetic conversations should not replace supervised clinical practice with real clients. The subtle fidelity violations we identified (Figure 1) would likely go unnoticed by non-specialist reviewers, highlighting the need for expert-guided evaluation of AI tools in clinical settings. Finally, our PE-specific metrics provide a framework for clinicians and developers to systematically evaluate AI-generated content for trauma treatment, potentially establishing minimum standards for therapeutic applications of synthetic data.
6.2 Supplementing Automated Metrics with Qualitative Fidelity Analysis
To supplement automated metrics, we conducted a manual review of synthetic dialogues, annotating exchanges for fidelity adherence or violations using Prolonged Exposure (PE) clinical guidelines. This qualitative analysis is essential, as statistical metrics alone may not detect subtle fidelity violations present even in linguistically fluent outputs. Fig. 1 presents five representative examples highlighting moments where such fluent outputs diverge from therapeutic fidelity, potentially undermining therapeutic efficacy. These include instances of role drift, premature processing, failure to maintain focus on exposure targets, improper SUDS implementation, and missed containment opportunities—violations that would often be imperceptible to both LLMs and lay reviewers. Fig. 1 illustrates these subtle violations clearly:
  • In Ex. 1, we observe role drift, where the therapist assumes the client’s narrative responsibility. While maintaining conversational coherence, this action undermines the therapeutic goal of client-driven processing.
  • Ex. 2 demonstrates a failure to maintain focus on exposure targets, allowing avoidance behavior that linguistic metrics cannot distinguish from therapeutically appropriate responses.
  • Ex. 3 and 4 show premature reflection and intervention, where the therapist interjects before full emotional processing occurs. This pattern can appear empathetic but critically disrupts the habituation processes central to PE efficacy.
  • Ex. 5 highlights improper SUDS implementation, which maintains the structural dialogue flow but compromises the validity of the clinical assessment.
These specific examples underscore why evaluation frameworks must integrate clinical expertise alongside computational metrics to accurately assess therapeutic fidelity in synthetic PE sessions.
7 Future Work
Future NLP research should focus on developing generative models with improved capabilities for tracking long-range dependencies and emotional dynamics (e.g., for Emotional Habituation, SUDS Progression), and creating automated metrics or classifiers trained with clinical expert annotations to better assess therapeutic process alignment.
8 Conclusions
This study evaluated synthetic therapeutic dialogues, showing that while they replicate structural features like turn-taking, they fall short in utterance length and conversational variability. Statistical analyses confirm these gaps, revealing that surface-level fluency can mask clinically meaningful fidelity lapses. We contribute a fidelity-aware evaluation framework tailored to Prolonged Exposure (PE) therapy, along with metrics highlighting that current generative models, despite their fluency, lack the nuance needed for high-stakes therapeutic contexts. Our findings underscore the risk of overestimating quality through non-expert or automated evaluations. For NLP, this work emphasizes the need for clinically grounded benchmarks, richer linguistic modeling, and methods like expert-in-the-loop validation and time-aware metrics to improve fidelity in domain-specific generation.
9 Ethical Concerns
All annotators involved in fidelity evaluation had prior clinical training or supervision, and data access was limited to IRB-approved investigators to ensure ethical compliance. The real-world dataset used in this study was collected under IRB-approved protocols with participant consent and cannot be shared publicly. For details on the ethical safeguards, simulation design, and usage guidelines related to the synthetic dataset, we refer readers to BN et al. (2025).
Limitations
While our study demonstrates the promise of synthetic dialogues in approximating real PE therapy interactions, several limitations remain. Although we measure fidelity using structural, statistical, and protocol-based metrics, our evaluation does not assess therapeutic effectiveness or downstream clinical outcomes. Synthetic data that aligns structurally with real dialogues may still fall short in supporting meaningful therapeutic engagement or behavior change, especially in high-stakes or emotionally complex scenarios. Moreover, we do not evaluate inter-rater agreement on fidelity violations, which is critical for establishing the robustness of manual annotations—this is planned for future work.
Appendix A Metric Definitions
We summarize some common PE terms  (Foa et al., 2007; Powers et al., 2010; Rauch et al., 2009; Hembree et al., 2003) below:
  1. 1.
    SUDS Progression: Change in Subjective Units of Distress Scale (SUDS) values mentioned by the client across the session.
  2. 2.
    Emotional Habituation: A decrease in distress levels or emotional intensity from the beginning to end of imaginal exposure.
  3. 3.
    Trauma Narrative Coherence: The syntactic and semantic coherence of the client’s trauma story, as measured by discourse coherence metrics.
  4. 4.
    Emotional Engagement: Captures the level of emotional expression, linked to better outcomes.
  5. 5.
    Avoidance Handling: Evaluates how well avoidance behaviors are addressed.
  6. 6.
    Exposure Guidance: Assesses the therapist’s role in structuring effective exposure exercises.
  7. 7.
    Cognitive Restructuring: Tracks how clients challenge maladaptive beliefs.
  8. 8.
    Avoidance Reduction: Quantifies improvements in engaging with trauma-related content.
  9. 9.
    Emotion Intensity: Assesses the variability and magnitude of emotional responses.
Appendix B Additional Tables
Table 2: Comparative analysis of real and synthetic data across multiple system-level metrics and their correlation values. The first two columns display the mean $\pm$ standard deviation for each metric, computed separately for real and synthetic datasets. The third and fourth columns provide the correlation values of these metrics within the real and synthetic datasets, respectively.
Metric Mean ± SD Correlation
Real Synth. Real Synth.
Norm. Spkr. Switches $0.99\pm 0.0$ $0.98\pm 0.0$ $0.85$ $0.72$
Norm. Total Turns $1.00\pm 0.0$ $1.00\pm 0.0$ $0.78$ $0.65$
Avg. Utt. Len. $68.7\pm 26.6$ $22.9\pm 1.7$ $0.91$ $0.76$
Utt. Length SD $135.9\pm 66.2$ $18.5\pm 2.3$ $0.89$ $0.74$
Norm. Avg. Turn Dur. $0.69\pm 0.6$ $0.12\pm 0.0$ $0.82$ $0.67$
Norm. Turn Dur. SD $1.38\pm 1.3$ $0.09\pm 0.0$ $0.79$ $0.64$
Norm. T Turns $0.50\pm 0.0$ $0.51\pm 0.0$ $0.87$ $0.71$
Norm. C Turns $0.50\pm 0.0$ $0.49\pm 0.0$ $0.86$ $0.70$
Norm. T Words $21.9\pm 13.2$ $4.9\pm 0.3$ $0.92$ $0.78$
Norm. C Words $46.8\pm 22.5$ $18.0\pm 1.7$ $0.90$ $0.75$
Turn Ratio (T/C) $0.01\pm 0.0$ $0.01\pm 0.0$ $0.80$ $0.66$
Word Ratio (T/C) $0.01\pm 0.0$ $0.00\pm 0$ 0.0 $0.82$ $0.68$
Vocab. Richness $0.13\pm 0.0$ $0.18\pm 0.0$ $0.77$ $0.63$
Readability Score $88.1\pm 4.6$ $89.2\pm 1.8$ $0.74$ $0.59$
Flow Entropy $1.30\pm 0.1$ $1.06\pm 0.0$ $0.88$ $0.73$
Avg. Perplexity $14.7\pm 2.3$ $21.2\pm 0.5$ $0.79$ $0.65$
Note: Norm. = Normalized, Utt. = Utterance, SD = Standard Deviation, Dur. = Duration, T = Therapist, C = Client.
Table 3: Mann-Whitney U test statistics and p-values comparing real and synthetic datasets.
Metric Statistic p-value
Norm. Speaker Switches $2.45\times 10^{3}$ ${\ll}0.001$
Norm. Total Turns $1.25\times 10^{3}$ $1.00$
Norm. Conv. Length $2.50\times 10^{3}$ ${\ll}0.001$
Avg. Utt. Length $2.50\times 10^{3}$ ${\ll}0.001$
Utt. Length Std $2.50\times 10^{3}$ ${\ll}0.001$
Norm. Turn Duration $2.49\times 10^{3}$ ${\ll}0.001$
Norm. Turn Dur. SD $2.50\times 10^{3}$ ${\ll}0.001$
Norm. T Turns $0.00$ ${\ll}0.001$
Norm. C Turns $2.50\times 10^{3}$ ${\ll}0.001$
Norm. T Words $2.50\times 10^{3}$ ${\ll}0.001$
Norm. C Words $2.47\times 10^{3}$ ${\ll}0.001$
T-C Turn Ratio $2.13\times 10^{3}$ $1.22\times 10^{-9}$
T-C Word Ratio $2.37\times 10^{3}$ ${\ll}0.001$
Vocabulary Richness $8.60\times 10^{1}$ ${\ll}0.001$
Readability Score $1.09\times 10^{3}$ $0.28$
Semantic Coherence $1.12\times 10^{3}$ $0.37$
Sem. Coherence Std $2.50\times 10^{3}$ ${\ll}0.001$
Flow Entropy $2.49\times 10^{3}$ ${\ll}0.001$
Avg. Perplexity $0.00$ ${\ll}0.001$
Local Coherence $1.12\times 10^{3}$ $0.37$
Coherence Std $2.50\times 10^{3}$ ${\ll}0.001$
Note: Norm. = Normalized, Conv. = Conversation, Utt. = Utterance, SD = Std. Dev, Dur. = Duration, T = Therapist, C = Client, Sem. = Semantic. $p\ll 0.001$ indicates $p<10^{-10}$.
Table 4: Relative importance of various conversational features in distinguishing real data from synthetic data, based on a predictive model. The Importance Score (%) reflects the contribution of each feature to the model’s decision-making process, with higher values indicating greater predictive power.
Feature Imp. Score (%)
Average Utterance Length $18.42$
Utterance Length Std Dev $15.76$
Normalized Therapist Words $12.58$
Normalized Client Words $10.94$
Flow Entropy $8.72$
Readability Score $7.89$
Normalized Avg Turn Duration $6.43$
Normalized Turn Duration Std $5.98$
Vocabulary Richness $4.85$
Average Perplexity $3.72$
Therapist-Client Turn Ratio $2.91$
Therapist-Client Word Ratio $2.32$
Normalized Speaker Switches $1.88$
Normalized Total Turns $0.98$
Table 5: Mann-Whitney U tests on PE therapy metrics comparing real and synthetic datasets.
Metric Statistic p-value
Trauma Narrative Coherence $2.31{\times}10^{3}$ ${\ll}0.001$
Emotional Engagement $2.45{\times}10^{3}$ ${\ll}0.001$
Avoidance Handling $1.98{\times}10^{3}$ $2.74{\times}10^{-7}$
Exposure Guidance $2.21{\times}10^{3}$ $5.62{\times}10^{-10}$
Cognitive Restructuring $2.12{\times}10^{3}$ $1.28{\times}10^{-8}$
Emotional Habituation $1.35{\times}10^{3}$ $0.102$
SUDS Progression $1.50{\times}10^{3}$ $0.073$
Avoidance Reduction $2.48{\times}10^{3}$ ${\ll}0.001$
Emotion Intensity $2.39{\times}10^{3}$ ${\ll}0.001$
Narrative Development $1.21{\times}10^{3}$ $0.251$

### v3

@@
 Abstract
-The growing adoption of synthetic data in healthcare is driven by privacy concerns, limited access to real-world data, and the high cost of annotation.
-This work explores the use of synthetic Prolonged Exposure (PE) therapeutic conversations for Post-Traumatic Stress Disorder (PTSD) as a scalable alternative for training and evaluating clinical models.
+The growing adoption of synthetic data in healthcare is driven by privacy concerns, limited access to real-world data, and high annotation costs.
+This work explores the use of synthetic Prolonged Exposure (PE) therapy conversations for Post-Traumatic Stress Disorder (PTSD) as a scalable alternative for training and evaluating clinical models.
 We systematically compare real and synthetic dialogues using linguistic, structural, and protocol-specific metrics, including turn-taking patterns and treatment fidelity.
-We also introduce and evaluate PE-specific metrics derived from linguistic analysis and semantic modeling, offering a novel framework for assessing clinical fidelity beyond surface fluency.
-Our findings show that although synthetic data holds promise for mitigating data scarcity and protecting patient privacy, it can struggle to capture the subtle dynamics of therapeutic interactions.
-In our dataset, synthetic dialogues match structural features of real-world dialogues (e.g., speaker switch ratio: 0.98 vs.
-0.99), however synthetic interactions do not adequately reflect key fidelity markers (e.g., distress monitoring).
-We highlight gaps in existing evaluation frameworks and advocate for fidelity-aware metrics that go beyond surface fluency to uncover clinically significant failures.
+We introduce and evaluate PE-specific metrics derived from linguistic analysis and semantic modeling, offering a novel framework for assessing clinical fidelity beyond surface fluency.
+Our findings show that while synthetic data holds promise for mitigating data scarcity and protecting patient privacy, it often struggles to capture the subtle dynamics of therapeutic interactions.
+Synthetic therapy dialogues closely match the structural features of real conversations (e.g., speaker switch ratio: 0.98 vs.
+0.99), but often fails to adequately reflect key fidelity markers such as distress monitoring.
+This work highlights gaps in current evaluation frameworks and advocate for fidelity-aware metrics that go beyond surface fluency to uncover clinically significant failures.
 Our findings clarify where synthetic data can effectively complement real-world datasets—and where critical limitations remain.
@@
 Evaluating Fidelity in Prolonged Exposure Dialogues
-Suhas BN1  Dominik Mattioli1  Saeed Abdullah1 Rosa I.
-Arriaga2  Chris W.
-Wiese3  Andrew M.
-Sherrill4 1College of Information Sciences and Technology, Penn State University, USA 2School of Interactive Computing, Georgia Tech, USA 3School of Psychology, Georgia Tech, USA 4Department of Psychiatry and Behavioral Sciences, Emory University, USA {bnsuhas,saeed}@psu.edu, andrew.m.sherrill@emory.edu
+Suhas BN1  Dominik Mattioli1  Andrew M.
+Sherrill2 Rosa I.
+Arriaga3  Chris W.
+Wiese4  Saeed Abdullah1 1College of Information Sciences and Technology, Penn State University, USA 2Department of Psychiatry and Behavioral Sciences, Emory University, USA 3School of Interactive Computing, Georgia Tech, USA 4School of Psychology, Georgia Tech, USA {bnsuhas,saeed}@psu.edu, andrew.m.sherrill@emory.edu
 1 Introduction
 Training machine learning models in sensitive domains like healthcare remains a challenge (Giuffrè and Shung, 2023; Kokosi and Harron, 2022).
-Real clinical conversations, particularly therapist-client dialogues, are rich sources of data for training models in areas such as mental health diagnostics, therapeutic interventions, and adherence monitoring.
-However, the high cost of data annotation, privacy concerns surrounding personally identifiable information (PII) (Kokosi and Harron, 2022; BN and Abdullah, 2022; BN et al., 2023), and ethical constraints in sharing sensitive interactions severely limit access to these datasets.
-Synthetic data has emerged as a promising alternative to address these challenges (Aher et al., 2023).
-By simulating conversations in a controlled environment, synthetic data offers scalability and privacy preservation while reducing the reliance on costly real-world annotations.
-In trauma-focused mental health care, especially Prolonged Exposure (PE) therapy for PTSD, large language models offer the potential to generate synthetic therapy dialogues at scale.
-Despite its potential, questions remain about the fidelity and utility of synthetic data to capture the subtle dynamics of therapeutic interactions (e.g., emotional depth, turn-taking patterns, and building therapeutic alliances) (Shen et al., 2024).
-Recent work (e.g., BN et al.
-(2025)) has shown that synthetic dialogues can fluently simulate PE sessions across diverse contexts.
-However, such outputs may only mimic surface structure and tone.
-Subtle fidelity lapses—like premature therapist reflection or avoidance reinforcement—often remain undetected by automatic metrics or non-clinical annotators (Chiu et al., 2024; Zhang et al., 2024; Lee et al., 2024b).
+Access to real clinical conversations—crucial for modeling tasks like mental health diagnostics and therapeutic dialogue understanding—is severely limited by high annotation costs, patient privacy concerns (Kokosi and Harron, 2022; BN and Abdullah, 2022; BN et al., 2023), and ethical constraints on data sharing.
+Synthetic data has emerged as a promising alternative, offering scalability and privacy preservation while reducing dependence on real-world annotations (Aher et al., 2023).
+In trauma-focused mental health care, particularly Prolonged Exposure (PE) therapy for PTSD, large language models (LLMs) can generate synthetic therapy dialogues at scale.
+However, questions remain about whether these dialogues capture more than surface-level fluency—specifically, the subtle dynamics of therapeutic fidelity such as emotional pacing, avoidance management, and protocol adherence (Shen et al., 2024).
+To address this, we develop and validate methods to measure clinically-relevant fidelity in generated dialogues—moving beyond standard metrics like coherence or perplexity.
+Prior work shows that while synthetic PE sessions can convincingly mimic real sessions in tone and structure (BN et al., 2025), they may still commit fidelity lapses, such as premature reflection or reinforcement of avoidance, which often go unnoticed by both automatic metrics and non-clinical annotators (Chiu et al., 2024; Zhang et al., 2024; Lee et al., 2024b).
 For instance, a therapist saying “That’s a really powerful insight” mid-exposure may appear empathic but violates PE protocol by derailing trauma processing (see Fig.
@@
 Without clinical expertise, both humans and automated metrics tend to overestimate fidelity.
-This paper introduces a fidelity-focused lens to evaluate whether synthetic PE sessions truly adhere to the therapeutic process, not just structure or tone.
-This fidelity-focused lens involves evaluating dialogues not only on linguistic coherence but also on adherence to specific therapeutic protocols and the successful navigation of key clinical interaction patterns (e.g., managing avoidance, appropriate SUDS monitoring) using both automated metrics and expert clinical assessment.
-In this study, we present a systematic comparison of real and synthetic clinical conversations, focusing on their linguistic, structural, and statistical characteristics.
-Our analysis explores whether synthetic data can replicate key features of real dialogues, including turn-taking patterns, protocol adherence, and emotional progression, while mitigating the risks associated with privacy and data scarcity.
-By analyzing both the strengths and limitations of synthetic therapy dialogues, our findings inform future improvements in the generation and evaluation of synthetic data for clinical applications.
-This study contributes to ongoing efforts to bridge the gap between real and synthetic data, providing a roadmap for the future of synthetic dialogue in sensitive domains.
-To our knowledge, this is the first study to combine statistical, linguistic, and clinical fidelity evaluations on both real and synthetic Prolonged Exposure (PE) therapy conversations at this scale.
+We introduce a fidelity-focused lens that evaluates dialogues on multiple dimensions: linguistic coherence, adherence to PE protocols, and the therapist’s navigation of key clinical interactions (e.g., managing avoidance, SUDS monitoring).
+This framework integrates automated scoring and expert clinical assessment for a more rigorous evaluation of synthetic dialogue quality.
+We present the first large-scale, multi-dimensional comparison of real and synthetic PE therapy sessions—analyzing their linguistic, structural, and clinical fidelity characteristics.
+Our findings demonstrate the current strengths and limitations of synthetic dialogues and inform future improvements in generation and evaluation methods.
+Our study contributes to bridging the real–synthetic data gap and provides a roadmap for advancing synthetic dialogue in sensitive healthcare domains.
 2 Motivation
-The development of synthetic clinical dialogues is motivated by several critical challenges in mental healthcare.
-Prolonged Exposure (PE) therapy—an evidence-based treatment for post-traumatic stress disorder (PTSD)—offers a structured therapeutic approach that could benefit from AI support, but faces significant obstacles:
-  1. 1.
-    Data scarcity: The limited availability of annotated clinical conversations restricts the development of AI systems for therapist training.
-Manual annotation is costly and time-consuming, often requiring up to a week per session per annotator, further compounding the scarcity of usable data.
-  2. 2.
-    Privacy constraints: Patient narratives contain highly sensitive personal information, making clinical data difficult to share or publish under privacy regulations and ethical guidelines.
-  3. 3.
-    Lack of diversity: Existing datasets often fail to capture diverse demographic profiles, trauma types, and therapeutic approaches, limiting model generalization.
-  4. 4.
-    Standardization challenges: In the absence of standardized benchmarks, fidelity evaluation of synthetic dialogues remains subjective and inconsistent.
-Our work addresses these challenges by creating and systematically evaluating synthetic PE therapy conversations that maintain clinical fidelity while protecting patient privacy.
-Beyond privacy preservation, synthetic data enables scalable annotation, broader demographic and trauma-type coverage, and controlled variation in dialogue structure—capabilities difficult to achieve with curated real-world data alone.
-Our framework compares real and synthetic dialogues across linguistic, structural, and clinical fidelity dimensions.
+Developing synthetic clinical dialogues is essential to overcome persistent barriers in mental health AI.
+In the context of Prolonged Exposure (PE) therapy, four challenges stand out: (1) data scarcity, due to the high cost and effort of annotation; (2) privacy constraints, which limit access to sensitive patient narratives; (3) lack of diversity, with datasets often failing to capture varied trauma types and demographics; and (4) evaluation inconsistency, as fidelity assessment lacks standardized benchmarks.
+This work directly addresses these issues by generating and evaluating synthetic PE dialogues that preserve protocol fidelity, enable scalable annotation, and support more equitable and robust training data for mental health applications.
 3 Related Work
-The evaluation of synthetic clinical conversation datasets, particularly for psychotherapy like Cognitive Behavioral Therapy (CBT) and general counseling, has grown significantly.
-However, research focusing on Prolonged Exposure (PE) therapy and other structured trauma-focused treatments remains limited.
-Existing studies primarily address (1) synthetic dataset generation and fidelity assessment, (2) evaluation metrics beyond lexical similarity, and (3) human-in-the-loop validation, often lacking PE-specific considerations.
+While synthetic clinical conversation datasets have advanced evaluations for general counseling and CBT, work specific to Prolonged Exposure (PE) and structured trauma-focused therapy is limited.
+Most studies focus on (1) synthetic dataset generation, (2) evaluation metrics beyond lexical similarity, and (3) human-in-the-loop validation, but generally lack PE-specific considerations.
 3.1 Synthetic Dataset Generation and Evaluation
-Several studies explore generating and assessing synthetic clinical datasets.
-For instance, Chiu et al.
-(2024) introduced BOLT to evaluate LLM-generated therapist behaviors (reflections, psychoeducation) in general counseling, while Qiu and Lan (2024) proposed SimPsyDial, benchmarking it against RealPsyDial using the Working Alliance Inventory (WAI-O-S).
-Zhang et al.
-(Zhang et al., 2024) presented CPsyCoun, reconstructing dialogues from reports and evaluating them with BERTScore (Zhang et al., 2019), GPTScore, and qualitative review.
-Others have focused on tailoring dialogues to specific counselor styles (Xie et al., 2024) or enhancing realism via data augmentation (Kim et al., 2024).
-Although datasets like Thousand Voices simulate structural variation (BN et al., 2025), current evaluations primarily compare synthetic data to real-world general counseling or CBT (Lee et al., 2024a; Shen et al., 2024).
-They generally lack explicit integration and assessment of trauma-specific markers or the structural fidelity crucial for PE, such as handling avoidance behaviors and managing imaginal exposure sequences.
-Our work extends these efforts by specifically evaluating whether synthetic PE sessions maintain clinical realism according to PE protocols (e.g., therapist containment, SUDS handling, avoidance redirection), highlighting fidelity violations where linguistic fluency masks procedural errors often missed by standard LLM evaluations or non-clinical reviewers.
+Recent work has explored synthetic clinical dialogue generation and assessment.
+BOLT evaluates LLM-generated therapist behaviors in general counseling (Chiu et al., 2024), while SimPsyDial benchmarks synthetic data using the Working Alliance Inventory (Qiu and Lan, 2024).
+CPsyCoun reconstructs dialogues for evaluation with BERTScore, GPTScore, and qualitative review (Zhang et al., 2024, 2019).
+Other studies focus on counselor style (Xie et al., 2024) and data augmentation (Kim et al., 2024).
+While Thousand Voices of Trauma introduces structural variation (BN et al., 2025), most evaluations compare to general counseling or CBT (Lee et al., 2024a; Shen et al., 2024) and rarely assess trauma-specific markers or PE’s structural fidelity, such as avoidance handling and imaginal exposure sequences.
+Our work directly addresses this by evaluating whether synthetic PE sessions capture clinical realism per PE protocols (e.g., SUDS, avoidance redirection), highlighting procedural errors missed by standard metrics.
 3.2 Beyond Lexical Metrics
-Evaluation is shifting from traditional lexical metrics (e.g., BLEU (Papineni et al., 2002)) towards more context-aware approaches.
-Transformer-based scores like BERTScore and GPTScore assess semantic alignment (Zhang et al., 2024; Xie et al., 2024), and psychological metrics (e.g., CTRS, PANAS in COUNSELINGEVAL (Lee et al., 2024a)) evaluate aspects like empathy and coherence.
-Specific interaction patterns, like Proactive Questioning Ability (PQA), have also been analyzed (Shen et al., 2024).
-However, these metrics are often unadapted for trauma-focused contexts, and advanced sequential techniques like Dynamic Time Warping (DTW), potentially useful for assessing turn-by-turn coherence in PE’s structured flow, remain largely unexplored in this domain.
-While qualitative assessments consider emotional tone and goal alignment (Zhang et al., 2024; Xie et al., 2024), and methods for structured questioning exist (Ren et al., 2024), robust quantitative frameworks for evaluating adherence to PE’s specific therapeutic structure are underdeveloped.
+Evaluation is shifting from lexical metrics like BLEU (Papineni et al., 2002) to semantic ones such as BERTScore and GPTScore (Zhang et al., 2024; Xie et al., 2024).
+Psychological measures (CTRS, PANAS) in COUNSELINGEVAL assess empathy and coherence (Lee et al., 2024a), and interactional features like PQA are explored (Shen et al., 2024).
+However, most metrics are not adapted for trauma therapy, and methods like Dynamic Time Warping (DTW)—potentially useful for structured PE flows—remain underused.
+While some qualitative reviews consider emotional tone and goal alignment (Zhang et al., 2024; Xie et al., 2024), and structured questioning is explored (Ren et al., 2024), robust frameworks for quantitative evaluation of PE-specific structure are lacking.
 3.3 Human-in-the-Loop Validation and Engagement Metrics
-Human expert validation is recognized as critical.
-Approaches like COUNSELINGEVAL (Lee et al., 2024a), PsyDT (Xie et al., 2024), and CPsyCoun (Zhang et al., 2024) incorporate clinician or professional reviewer feedback to assess quality, empathy, and safety.
-Concurrently, personalization and engagement metrics are emphasized (Lee et al., 2024a; Zhang et al., 2024; Xie et al., 2024), and prior work addresses empathy (Morris et al., 2018) and structured workflows (Ren et al., 2024).
-However, these human evaluations predominantly focus on general counseling or CBT contexts.
-They typically omit systematic assessment of PE-specific elements like trauma cue processing, therapeutic handling of emotional regulation, and adherence to PE protocols, especially when compared against real-world PE therapy data.
+Clinician validation is now common, as in COUNSELINGEVAL (Lee et al., 2024a), PsyDT (Xie et al., 2024), and CPsyCoun (Zhang et al., 2024), which assess quality, empathy, and safety.
+Metrics for engagement and personalization are also used (Lee et al., 2024a; Zhang et al., 2024; Xie et al., 2024).
+Prior work covers empathy (Morris et al., 2018) and structured workflows (Ren et al., 2024).
+However, human evaluation remains focused on general counseling or CBT, rarely addressing PE-specific components like trauma cue processing or protocol adherence compared to real PE data.
 3.4 Bridging the Gap in PE Therapy Evaluation
-Prolonged Exposure (PE) therapy is a first-line, empirically supported treatment for PTSD (Sherrill and Rauch, 2019).
-Despite progress in synthetic dialogue generation (Qiu and Lan, 2024; Zhang et al., 2024; Xie et al., 2024), evaluation metrics (Lee et al., 2024a; Zhang et al., 2024; Xie et al., 2024), and human validation (Lee et al., 2024a; Zhang et al., 2024; Xie et al., 2024), a significant gap persists in assessing the fidelity of synthetic data for structured, trauma-focused therapies like PE (Chiu et al., 2024; Qiu and Lan, 2024; Lee et al., 2024a; Shen et al., 2024; Zhang et al., 2024; Xie et al., 2024; Hu et al., 2024).
-As discussed, current approaches generally lack the specific focus and metrics needed to evaluate PE’s unique conversational structures (e.g., avoidance handling, trauma cue processing, imaginal exposure flow) and robustness to therapy-specific challenges like patient emotional dysregulation or engagement difficulties (Qiu and Lan, 2024; Lee et al., 2024a; Shen et al., 2024; Zhang et al., 2024; Xie et al., 2024; Kim et al., 2024).
-Our work directly addresses this limitation by systematically evaluating synthetic data alignment with structured PE interactions, identifying key areas for improvement in generating clinically valid PE therapy dialogues (See Table 1 and Appendix A).
+PE is a first-line treatment for PTSD (Sherrill and Rauch, 2019; Varkovitzky et al., 2018; Rauch et al., 2021; Ragsdale et al., 2020; Evans et al., 2020; Yasinski et al., 2017).
+Despite progress in synthetic dialogue generation (Qiu and Lan, 2024; Zhang et al., 2024; Xie et al., 2024), evaluation metrics (Lee et al., 2024a; Zhang et al., 2024; Xie et al., 2024), and human validation, a gap remains in fidelity assessment for trauma-focused therapies like PE (Chiu et al., 2024; Qiu and Lan, 2024; Lee et al., 2024a; Shen et al., 2024; Zhang et al., 2024; Xie et al., 2024; Hu et al., 2024).
+Current approaches lack metrics and focus tailored to PE’s unique conversational structure—including avoidance handling, trauma cue processing, and imaginal exposure flow—and do not address challenges such as patient emotional dysregulation (Qiu and Lan, 2024; Lee et al., 2024a; Shen et al., 2024; Zhang et al., 2024; Xie et al., 2024; Kim et al., 2024).
+Our work systematically evaluates synthetic data for alignment with PE structure and clinical validity, identifying areas for improving synthetic PE dialogue quality (see Table 1, Appendix A).
 Metric Yes No N/A
@@
     Linguistic complexity and coherence: Average Utterance Length, Utterance Length Std Dev, and Readability Score assess how natural and readable the synthetic text is.
-A significant deviation from real data would indicate a lack of coherence in generated responses.
+Significant deviations indicate poor coherence.
   3. 3.
@@
 While mean comparisons provide a general overview, correlation analysis (see Table 3) quantifies the consistency of relationships across metrics between real and synthetic data.
-This step helps us determine whether patterns observed in real data persist in synthetic dialogues.
+This determines if real-data patterns persist in synthetic dialogues.
 For instance:
@@
 To confirm whether differences between real and synthetic data are statistically meaningful, we conduct Mann-Whitney U tests (Table 4).
-This analysis assesses whether observed variations are due to randomness or inherent inconsistencies in the synthetic data.
+This assesses if variations are random or due to inherent inconsistencies.
 Key takeaways include:
@@
 4.1.4 PE-Specific Metrics
-Prolonged Exposure (PE) therapy reduces pathological fear through repeated trauma exposure.
+PE therapy reduces pathological fear through repeated trauma exposure.
 To assess the fidelity of synthetic PE sessions, we evaluate key therapeutic constructs (See Appendix A for definitions and Table 5): Trauma Narrative Coherence measures how structured and detailed a client’s trauma account is, reflecting cognitive integration.
@@
 These metrics are derived using linguistic analysis, semantic modeling, and interaction patterns.
+4.1.5 Qualitative Fidelity Assessment
+To supplement automated evaluation metrics, we conducted a manual review of synthetic dialogues, annotating exchanges for fidelity adherence or violations using established PE clinical guidelines.
+Manual fidelity annotations were then reviewed by a licensed clinical psychotherapist for adherence to PE protocol.
+While inter-rater agreement was not computed in this exploratory phase, these annotations served as a qualitative tool to identify and illustrate typical fidelity lapses—especially those potentially overlooked by automated metrics, rather than as quantitative ground truth.
+Representative examples of such lapses identified through this process are illustrated in Figure 1 and discussed in Section 6.2.
 5 Findings
@@
 The system-level metrics comparison highlights alignment between real and synthetic dialogues, revealing both structural matches and model-driven differences.
-While turn-taking patterns (e.g., Normalized Speaker Switches, Therapist-Client Turn Ratio, and Normalized Therapist/Client Turns) remain similar, variations in utterance length, lexical diversity, and entropy-based measures arise due to prolonged exposure therapy and model constraints.
+While turn-taking patterns (e.g., Normalized Speaker Switches, Therapist-Client Turn Ratio, and Normalized Therapist/Client Turns) remain similar, variations in utterance length, lexical diversity, and entropy-based measures arise due to PE therapy and model constraints.
   1. 1.
@@
 Lower correlations in linguistic complexity metrics (e.g., Readability Score, Vocabulary Richness, Perplexity) reflect structural differences but do not hinder the coherence/naturalness of the conversation.
-5.2 Statistical Significance Testing
-The Mann-Whitney U test results confirm that many observed differences between real and synthetic dialogues are statistically significant ($p<0.05$).
-However, these differences stem from model design choices and practical constraints rather than fundamental shortcomings.
-Key observations include:
-  1. 1.
-    Distinct patterns in utterance structure: The test shows differences in utterance length, turn duration, and their standard deviations ($p<10^{-17}$).
-This is expected, as synthetic dialogues are designed to maintain coherence by producing more structured and concise responses.
-In real PE therapy, clients occasionally have extended monologues, which LLMs struggle to handle due to context window limitations.
-To compensate, we shorten utterances while preserving the conversational structure.
-Methods like Chain-of-Thought prompting have improved this, but achieving full parity would require larger output contexts.
-  2. 2.
-    Lexical properties follow a structured pattern: Differences in Vocabulary Richness ($p<10^{-15}$) and Flow Entropy ($p<10^{-17}$) indicate that synthetic data exhibits a more varied vocabulary but within a constrained framework.
-This is a direct result of the model prioritizing coherence and avoiding redundant expressions.
-While real conversations naturally contain more spontaneity, the synthetic approach ensures stability in generated dialogue while maintaining conversational depth.
-  3. 3.
-    Certain aspects remain comparable: Metrics such as Normalized Total Turns ($p=1.00$) and Readability Score ($p=0.28$) show no significant differences, meaning that despite shorter utterances, the number of conversational exchanges and overall readability remain aligned with real data.
-This suggests that while individual responses may be more concise, the overall flow and engagement in the conversation are well-preserved.
-These findings reinforce that the synthetic model captures key conversational characteristics while ensuring structured, coherent responses.
-While differences exist, they align with known model constraints and do not compromise the overall integrity of the generated dialogues.
-5.3 Feature Importance
-Feature importance analysis highlights which metrics most influence synthetic data generation, reinforcing trends observed in prior evaluations.
-While not the primary focus, these insights provide valuable confirmation of the model’s alignment with conversational structure and linguistic patterns.
-  1. 1.
-    Turn-taking features contribute heavily: Metrics such as Normalized Speaker Switches and Therapist-Client Turn Ratio emerge as highly influential.
-This is a positive indication that the model effectively captures conversational flow, ensuring balanced exchanges between therapist and client.
-The strong correlation between real and synthetic data in these aspects suggests that the generated dialogues maintain natural interaction patterns.
-  2. 2.
-    Utterance length and entropy-based measures shape responses: Metrics like Average Utterance Length, Perplexity, and Flow Entropy stand out as key distinguishing factors.
-This aligns with statistical significance results, reflecting that synthetic responses are more structured and predictable.
-This behavior is expected, as the model prioritizes coherence and consistency, which helps maintain clarity even in complex dialogues.
-While synthetic responses may be more concise, they still reflect key interaction patterns but require further refinement to support emotionally dynamic exchanges.
-  3. 3.
-    Lexical richness plays a secondary role: Vocabulary Richness was moderately influential, indicating that synthetic data maintains a varied vocabulary while ensuring structured dialogue.
-While real-world conversations may exhibit greater linguistic flexibility, synthetic dialogues remain clear and contextually appropriate, balancing variation with coherence.
-5.4 Prolonged Exposure Specific Metrics
-Table 5 presents the statistical test results comparing synthetic Prolonged Exposure (PE) data with real therapy session data.
-The results indicate that the synthetic data effectively captures key therapeutic constructs, with several core PE metrics such as Trauma Narrative Coherence, Emotional Engagement, Avoidance Handling, Exposure Guidance, Cognitive Restructuring, Avoidance Reduction, and Emotion Intensity showing strong alignment with real data (p < 0.001) showing strong alignment with real data.
-These results suggest that the synthetic data successfully replicate the structure and engagement patterns seen in real therapy sessions.
-However, certain metrics do not reach statistical significance, indicating areas where synthetic data generation may need refinement.
-Emotional Habituation ($p=0.102$) and SUDS Progression ($p=0.073$) track distress reduction over repeated exposures, and their lack of significance suggests the synthetic data may not fully model this progressive decrease.
-Similarly, while Trauma Narrative Coherence is well-captured, Narrative Development ($p=0.251$) appears to be less robust, possibly due to limited context retention or rigid structure in generated responses.
-Despite these discrepancies, the overall performance of synthetic data suggests it is a viable alternative for privacy-sensitive applications.
-Addressing dynamic narrative evolution could further improve alignment with real data.
-However, fidelity violations in synthetic sessions are often subtle and may not disrupt structural metrics.
-For example, therapist utterances can appear empathetic or affirming while inadvertently shifting the session away from trauma anchoring.
-These moments often go unnoticed by automated scoring or non-clinical reviewers, underscoring the need for fidelity-aware evaluation frameworks that integrate human clinical judgment.
-6 Discussion
-In this paper, we evaluate the fidelity of synthetic Prolonged Exposure (PE) therapy conversations by comparing them to real interactions using linguistic, structural, and statistical analyses.
-We discuss the findings from four key perspectives: system-level metrics and correlation (Table 2), statistical significance testing (Table 3), feature importance (Table 4), and PE-specific metrics (Table 5).
-6.1 Clinical Implications
-Our findings have several important implications for clinical practice.
-First, synthetic PE sessions could serve as valuable training tools for novice therapists learning to identify protocol deviations, as they provide controlled examples of both adherent and non-adherent interactions without privacy concerns.
-Second, the consistent replication of structural elements (turn-taking, session flow) suggests synthetic data can supplement clinical training materials, particularly in regions with limited access to specialized trauma training.
-However, the identified limitations in capturing emotional habituation patterns and dynamic narrative development indicate synthetic conversations should not replace supervised clinical practice with real clients.
-The subtle fidelity violations we identified (Figure 1) would likely go unnoticed by non-specialist reviewers, highlighting the need for expert-guided evaluation of AI tools in clinical settings.
-Finally, our PE-specific metrics provide a framework for clinicians and developers to systematically evaluate AI-generated content for trauma treatment, potentially establishing minimum standards for therapeutic applications of synthetic data.
-6.2 Supplementing Automated Metrics with Qualitative Fidelity Analysis
-To supplement automated metrics, we conducted a manual review of synthetic dialogues, annotating exchanges for fidelity adherence or violations using Prolonged Exposure (PE) clinical guidelines.
-This qualitative analysis is essential, as statistical metrics alone may not detect subtle fidelity violations present even in linguistically fluent outputs.
-Fig.
-1 presents five representative examples highlighting moments where such fluent outputs diverge from therapeutic fidelity, potentially undermining therapeutic efficacy.
-These include instances of role drift, premature processing, failure to maintain focus on exposure targets, improper SUDS implementation, and missed containment opportunities—violations that would often be imperceptible to both LLMs and lay reviewers.
-Fig.
-1 illustrates these subtle violations clearly:
-  • In Ex.
-1, we observe role drift, where the therapist assumes the client’s narrative responsibility.
-While maintaining conversational coherence, this action undermines the therapeutic goal of client-driven processing.
-  • Ex. 2 demonstrates a failure to maintain focus on exposure targets, allowing avoidance behavior that linguistic metrics cannot distinguish from therapeutically appropriate responses.
-  • Ex.
-3 and 4 show premature reflection and intervention, where the therapist interjects before full emotional processing occurs.
-This pattern can appear empathetic but critically disrupts the habituation processes central to PE efficacy.
-  • Ex. 5 highlights improper SUDS implementation, which maintains the structural dialogue flow but compromises the validity of the clinical assessment.
-These specific examples underscore why evaluation frameworks must integrate clinical expertise alongside computational metrics to accurately assess therapeutic fidelity in synthetic PE sessions.
-7 Future Work
-Future NLP research should focus on developing generative models with improved capabilities for tracking long-range dependencies and emotional dynamics (e.g., for Emotional Habituation, SUDS Progression), and creating automated metrics or classifiers trained with clinical expert annotations to better assess therapeutic process alignment.
-8 Conclusions
-This study evaluated synthetic therapeutic dialogues, showing that while they replicate structural features like turn-taking, they fall short in utterance length and conversational variability.
-Statistical analyses confirm these gaps, revealing that surface-level fluency can mask clinically meaningful fidelity lapses.
-We contribute a fidelity-aware evaluation framework tailored to Prolonged Exposure (PE) therapy, along with metrics highlighting that current generative models, despite their fluency, lack the nuance needed for high-stakes therapeutic contexts.
-Our findings underscore the risk of overestimating quality through non-expert or automated evaluations.
-For NLP, this work emphasizes the need for clinically grounded benchmarks, richer linguistic modeling, and methods like expert-in-the-loop validation and time-aware metrics to improve fidelity in domain-specific generation.
-9 Ethical Concerns
-All annotators involved in fidelity evaluation had prior clinical training or supervision, and data access was limited to IRB-approved investigators to ensure ethical compliance.
-The real-world dataset used in this study was collected under IRB-approved protocols with participant consent and cannot be shared publicly.
-For details on the ethical safeguards, simulation design, and usage guidelines related to the synthetic dataset, we refer readers to BN et al.
-(2025).
-Limitations
-While our study demonstrates the promise of synthetic dialogues in approximating real PE therapy interactions, several limitations remain.
-Although we measure fidelity using structural, statistical, and protocol-based metrics, our evaluation does not assess therapeutic effectiveness or downstream clinical outcomes.
-Synthetic data that aligns structurally with real dialogues may still fall short in supporting meaningful therapeutic engagement or behavior change, especially in high-stakes or emotionally complex scenarios.
-Moreover, we do not evaluate inter-rater agreement on fidelity violations, which is critical for establishing the robustness of manual annotations—this is planned for future work.
-Appendix A Metric Definitions
-We summarize some common PE terms  (Foa et al., 2007; Powers et al., 2010; Rauch et al., 2009; Hembree et al., 2003) below:
-  1. 1.
-    SUDS Progression: Change in Subjective Units of Distress Scale (SUDS) values mentioned by the client across the session.
-  2. 2.
-    Emotional Habituation: A decrease in distress levels or emotional intensity from the beginning to end of imaginal exposure.
-  3. 3.
-    Trauma Narrative Coherence: The syntactic and semantic coherence of the client’s trauma story, as measured by discourse coherence metrics.
-  4. 4.
-    Emotional Engagement: Captures the level of emotional expression, linked to better outcomes.
-  5. 5.
-    Avoidance Handling: Evaluates how well avoidance behaviors are addressed.
-  6. 6.
-    Exposure Guidance: Assesses the therapist’s role in structuring effective exposure exercises.
-  7. 7.
-    Cognitive Restructuring: Tracks how clients challenge maladaptive beliefs.
-  8. 8.
-    Avoidance Reduction: Quantifies improvements in engaging with trauma-related content.
-  9. 9.
-    Emotion Intensity: Assesses the variability and magnitude of emotional responses.
-Appendix B Additional Tables
 Table 2: Comparative analysis of real and synthetic data across multiple system-level metrics and their correlation values.
@@
 Note: Norm. = Normalized, Utt. = Utterance, SD = Standard Deviation, Dur. = Duration, T = Therapist, C = Client.
+5.2 Statistical Significance Testing
+The Mann-Whitney U test results confirm that many observed differences between real and synthetic dialogues are statistically significant ($p<0.05$).
+However, these differences stem from model design choices and practical constraints rather than fundamental shortcomings.
+Key observations include:
+  1. 1.
+    Distinct patterns in utterance structure: The test shows differences in utterance length, turn duration, and their standard deviations ($p<10^{-17}$).
+This is expected, as synthetic dialogues are designed to maintain coherence by producing more structured and concise responses.
+In real PE therapy, clients occasionally have extended monologues, which LLMs struggle to handle due to context window limitations.
+To compensate, we shorten utterances while preserving the conversational structure.
+Methods like Chain-of-Thought prompting have improved this, but achieving full parity would require larger output contexts.
+  2. 2.
+    Lexical properties follow a structured pattern: Differences in Vocabulary Richness ($p<10^{-15}$) and Flow Entropy ($p<10^{-17}$) indicate that synthetic data exhibits a more varied vocabulary but within a constrained framework.
+This is a direct result of the model prioritizing coherence and avoiding redundant expressions.
+While real conversations naturally contain more spontaneity, the synthetic approach ensures stability in generated dialogue while maintaining conversational depth.
+  3. 3.
+    Certain aspects remain comparable: Metrics such as Normalized Total Turns ($p=1.00$) and Readability Score ($p=0.28$) show no significant differences, meaning that despite shorter utterances, the number of conversational exchanges and overall readability remain aligned with real data.
+This suggests that while individual responses may be more concise, the overall flow and engagement in the conversation are well-preserved.
+These findings reinforce that the synthetic model captures key conversational characteristics while ensuring structured, coherent responses.
+While differences exist, they align with known model constraints and do not compromise the overall integrity of the generated dialogues.
 Table 3: Mann-Whitney U test statistics and p-values comparing real and synthetic datasets.
 Metric Statistic p-value
-Norm. Speaker Switches $2.45\times 10^{3}$ ${\ll}0.001$
+Norm. Speaker Switches $2.45\times 10^{3}$ $p<0.001$
 Norm. Total Turns $1.25\times 10^{3}$ $1.00$
-Norm. Conv. Length $2.50\times 10^{3}$ ${\ll}0.001$
-Avg. Utt. Length $2.50\times 10^{3}$ ${\ll}0.001$
-Utt. Length Std $2.50\times 10^{3}$ ${\ll}0.001$
-Norm. Turn Duration $2.49\times 10^{3}$ ${\ll}0.001$
-Norm. Turn Dur. SD $2.50\times 10^{3}$ ${\ll}0.001$
-Norm. T Turns $0.00$ ${\ll}0.001$
-Norm. C Turns $2.50\times 10^{3}$ ${\ll}0.001$
-Norm. T Words $2.50\times 10^{3}$ ${\ll}0.001$
-Norm. C Words $2.47\times 10^{3}$ ${\ll}0.001$
+Norm. Conv. Length $2.50\times 10^{3}$ $p<0.001$
+Avg. Utt. Length $2.50\times 10^{3}$ $p<0.001$
+Utt. Length Std $2.50\times 10^{3}$ $p<0.001$
+Norm. Turn Duration $2.49\times 10^{3}$ $p<0.001$
+Norm. Turn Dur. SD $2.50\times 10^{3}$ $p<0.001$
+Norm. T Turns $0.00$ $p<0.001$
+Norm. C Turns $2.50\times 10^{3}$ $p<0.001$
+Norm. T Words $2.50\times 10^{3}$ $p<0.001$
+Norm. C Words $2.47\times 10^{3}$ $p<0.001$
 T-C Turn Ratio $2.13\times 10^{3}$ $1.22\times 10^{-9}$
-T-C Word Ratio $2.37\times 10^{3}$ ${\ll}0.001$
-Vocabulary Richness $8.60\times 10^{1}$ ${\ll}0.001$
+T-C Word Ratio $2.37\times 10^{3}$ $p<0.001$
+Vocabulary Richness $8.60\times 10^{1}$ $p<0.001$
 Readability Score $1.09\times 10^{3}$ $0.28$
 Semantic Coherence $1.12\times 10^{3}$ $0.37$
-Sem. Coherence Std $2.50\times 10^{3}$ ${\ll}0.001$
-Flow Entropy $2.49\times 10^{3}$ ${\ll}0.001$
-Avg. Perplexity $0.00$ ${\ll}0.001$
+Sem. Coherence Std $2.50\times 10^{3}$ $p<0.001$
+Flow Entropy $2.49\times 10^{3}$ $p<0.001$
+Avg. Perplexity $0.00$ $p<0.001$
 Local Coherence $1.12\times 10^{3}$ $0.37$
-Coherence Std $2.50\times 10^{3}$ ${\ll}0.001$
-Note: Norm. = Normalized, Conv. = Conversation, Utt. = Utterance, SD = Std. Dev, Dur. = Duration, T = Therapist, C = Client, Sem. = Semantic. $p\ll 0.001$ indicates $p<10^{-10}$.
+Coherence Std $2.50\times 10^{3}$ $p<0.001$
+Note: Norm. = Normalized, Conv. = Conversation, Utt. = Utterance, SD = Std. Dev, Dur. = Duration, T = Therapist, C = Client, Sem. = Semantic. $p<0.001$ indicates $p<10^{-10}$.
+5.3 Feature Importance
+Feature importance analysis identifies the metrics most influencing synthetic data generation, confirming trends from previous evaluations.
+While not the central focus, these results support the model’s alignment with conversational structure and linguistic patterns.
+  1. 1.
+    Turn-taking features are most influential: Metrics such as Normalized Speaker Switches and Therapist-Client Turn Ratio dominate, indicating the model effectively captures conversational flow and balanced exchanges.
+Strong correlations with real data suggest natural interaction patterns are preserved.
+  2. 2.
+    Utterance length and entropy shape responses: Average Utterance Length, Perplexity, and Flow Entropy are key distinguishing factors.
+Synthetic responses are more structured and predictable, prioritizing coherence and consistency.
+While often more concise, they capture essential interaction patterns, though further refinement is needed for richer emotional dynamics.
+  3. 3.
+    Lexical richness is moderately important: Vocabulary Richness plays a secondary role, reflecting that synthetic dialogues maintain varied but structured language.
+While real conversations are more flexible, synthetic responses balance vocabulary diversity with clarity.
+5.4 Prolonged Exposure Specific Metrics
+Table 5 presents the statistical test results comparing synthetic Prolonged Exposure data with real therapy session data.
+The results indicate that the synthetic data effectively captures key therapeutic constructs, with several core PE metrics such as Trauma Narrative Coherence, Emotional Engagement, Avoidance Handling, Exposure Guidance, Cognitive Restructuring, Avoidance Reduction, and Emotion Intensity showing strong alignment with real data (p < 0.001) showing strong alignment with real data.
+These results suggest that the synthetic data successfully replicate the structure and engagement patterns seen in real therapy sessions.
+However, certain metrics do not reach statistical significance, indicating areas where synthetic data generation may need refinement.
+Emotional Habituation ($p=0.102$) and SUDS Progression ($p=0.073$) track distress reduction over repeated exposures, and their lack of significance suggests the synthetic data may not fully model this progressive decrease.
+Similarly, while Trauma Narrative Coherence is well-captured, Narrative Development ($p=0.251$) appears to be less robust, possibly due to limited context retention or rigid structure in generated responses.
+Despite these discrepancies, the overall performance of synthetic data suggests it is a viable alternative for privacy-sensitive applications.
+Addressing dynamic narrative evolution could further improve alignment with real data.
+However, fidelity violations in synthetic sessions are often subtle and may not disrupt structural metrics.
+For example, therapist utterances can appear empathetic or affirming while inadvertently shifting the session away from trauma anchoring.
+These moments often go unnoticed by automated scoring or non-clinical reviewers, underscoring the need for fidelity-aware evaluation frameworks that integrate human clinical judgment.
+6 Discussion
+This paper evaluated the fidelity of synthetic PE therapy conversations by comparing them to real interactions using linguistic, structural, and statistical analyses.
+We discuss the findings from four key perspectives: system-level metrics and correlation (Table 2), statistical significance testing (Table 3), feature importance (Table 4), and PE-specific metrics (Table 5).
+6.1 Clinical Implications
+Our findings have several important implications for clinical practice.
+First, synthetic PE sessions can serve as valuable training tools for novice therapists learning to identify protocol deviations, as they offer controlled examples of both adherent and non-adherent interactions without privacy concerns.
+Second, the consistent replication of structural elements (e.g., turn-taking, session flow) indicates that synthetic data can effectively supplement clinical training materials, particularly in settings with limited access to specialized trauma training.
+However, our qualitative analysis reveals limitations that automated metrics may overlook.
+As shown in Figure 1, even linguistically fluent synthetic dialogues can contain subtle but clinically significant fidelity violations—including role drift, premature processing, and improper SUDS implementation—that typically escape quantitative detection.
+These lapses would likely go unnoticed by non-specialist reviewers, highlighting the necessity of expert-guided evaluation for AI tools in clinical contexts.
+Finally, our PE-specific metrics provide a systematic framework for clinicians and developers to assess AI-generated content for trauma treatment, potentially setting minimum standards for therapeutic applications of synthetic data.
+6.2 Qualitative Analysis of Fidelity Lapses
+Section 6.1 highlighted the clinical implications of subtle fidelity violations that often escape automated detection.
+To provide a more granular understanding of these critical lapses and illustrate the necessity of expert review, Figure 1 details five representative examples identified through our qualitative analysis.
+These specific instances demonstrate common patterns of deviation from PE protocol:
+  • Ex. 1: Role Drift — The therapist assumes the client’s narrative responsibility, undermining the therapeutic goal of client-driven processing.
+  • Ex. 2: Failure to Maintain Focus on Exposure Targets — Avoidance behavior is permitted, a lapse not distinguishable by standard metrics.
+  • Ex. 3–4: Premature Reflection and Intervention — The therapist interjects before full emotional processing, disrupting the habituation process crucial for PE efficacy.
+  • Ex. 5: Improper SUDS Implementation — The dialogue maintains structural flow, but the clinical assessment validity is compromised.
+These detailed examples underscore why evaluation frameworks must integrate clinical expertise alongside computational metrics to accurately assess therapeutic fidelity in synthetic PE sessions.
+7 Future Work
+Future NLP research should focus on developing generative models with improved capabilities for tracking long-range dependencies and emotional dynamics (e.g., for Emotional Habituation, SUDS Progression), and creating automated metrics or classifiers trained with clinical expert annotations to better assess therapeutic process alignment.
+8 Conclusions
+This study evaluated synthetic therapeutic dialogues, showing that while they replicate structural features like turn-taking, they fall short in utterance length and conversational variability.
+Statistical analyses confirm these gaps, revealing that surface-level fluency can mask clinically meaningful fidelity lapses.
+We contribute a fidelity-aware evaluation framework tailored to Prolonged Exposure (PE) therapy, along with metrics highlighting that current generative models, despite their fluency, lack the nuance needed for high-stakes therapeutic contexts.
+Our findings underscore the risk of overestimating quality through non-expert or automated evaluations.
+For NLP, this work emphasizes the need for clinically grounded benchmarks, richer linguistic modeling, and methods like expert-in-the-loop validation and time-aware metrics to improve fidelity in domain-specific generation.
+9 Ethical Concerns
+All annotators involved in fidelity evaluation had prior clinical training or supervision, and data access was limited to IRB-approved investigators to ensure ethical compliance.
+The real-world dataset used in this study was collected under IRB-approved protocols with participant consent and cannot be shared publicly.
+For details on the ethical safeguards, simulation design, and usage guidelines related to the synthetic dataset, we refer readers to BN et al.
+(2025).
+Limitations
+While our study demonstrates the promise of synthetic dialogues in approximating real PE therapy interactions, several limitations remain.
+Although we measure fidelity using structural, statistical, and protocol-based metrics, our evaluation does not assess therapeutic effectiveness or downstream clinical outcomes.
+Synthetic data that aligns structurally with real dialogues may still fall short in supporting meaningful therapeutic engagement or behavior change, especially in high-stakes or emotionally complex scenarios.
+Moreover, we do not evaluate inter-rater agreement on fidelity violations, which is critical for establishing the robustness of manual annotations—this is planned for future work.
+Acknowledgement
+This work is supported by the National Science Foundation (NSF) under Grant No.
+(Details omitted for double-blind review.) Any opinions, findings, & conclusions or recommendations expressed in this work are those of the author(s) and do not necessarily reflect the views of the NSF.
+Appendix A Metric Definitions
+Key PE terms (Foa et al., 2007; Powers et al., 2010; Rauch et al., 2009; Hembree et al., 2003):
+  1. 1.
+    SUDS Progression: Change in Subjective Units of Distress (SUDS) reported by the client across the session.
+  2. 2.
+    Emotional Habituation: Decrease in distress or emotional intensity from start to end of imaginal exposure.
+  3. 3.
+    Trauma Narrative Coherence: Syntactic and semantic coherence of the trauma narrative, via discourse metrics.
+  4. 4.
+    Emotional Engagement: Degree of emotional expression, associated with better outcomes.
+  5. 5.
+    Avoidance Handling: Effectiveness in addressing avoidance behaviors.
+  6. 6.
+    Exposure Guidance: Therapist’s structuring of exposure exercises.
+  7. 7.
+    Cognitive Restructuring: Client’s efforts to challenge maladaptive beliefs.
+  8. 8.
+    Avoidance Reduction: Increased engagement with trauma content.
+  9. 9.
+    Emotion Intensity: Variability and magnitude of emotional responses.
+Appendix B Additional Tables
 Table 4: Relative importance of various conversational features in distinguishing real data from synthetic data, based on a predictive model.
@@
 Metric Statistic p-value
-Trauma Narrative Coherence $2.31{\times}10^{3}$ ${\ll}0.001$
-Emotional Engagement $2.45{\times}10^{3}$ ${\ll}0.001$
+Trauma Narrative Coherence $2.31{\times}10^{3}$ $p<0.001$
+Emotional Engagement $2.45{\times}10^{3}$ $p<0.001$
 Avoidance Handling $1.98{\times}10^{3}$ $2.74{\times}10^{-7}$
@@
 SUDS Progression $1.50{\times}10^{3}$ $0.073$
-Avoidance Reduction $2.48{\times}10^{3}$ ${\ll}0.001$
-Emotion Intensity $2.39{\times}10^{3}$ ${\ll}0.001$
+Avoidance Reduction $2.48{\times}10^{3}$ $p<0.001$
+Emotion Intensity $2.39{\times}10^{3}$ $p<0.001$
 Narrative Development $1.21{\times}10^{3}$ $0.251$
+SUDS = Subjective Units of Distress Scale. $p<0.001$ indicates $p<10^{-10}$.

### v4

@@
 Abstract
-The growing adoption of synthetic data in healthcare is driven by privacy concerns, limited access to real-world data, and high annotation costs.
-This work explores the use of synthetic Prolonged Exposure (PE) therapy conversations for Post-Traumatic Stress Disorder (PTSD) as a scalable alternative for training and evaluating clinical models.
-We systematically compare real and synthetic dialogues using linguistic, structural, and protocol-specific metrics, including turn-taking patterns and treatment fidelity.
-We introduce and evaluate PE-specific metrics derived from linguistic analysis and semantic modeling, offering a novel framework for assessing clinical fidelity beyond surface fluency.
-Our findings show that while synthetic data holds promise for mitigating data scarcity and protecting patient privacy, it often struggles to capture the subtle dynamics of therapeutic interactions.
-Synthetic therapy dialogues closely match the structural features of real conversations (e.g., speaker switch ratio: 0.98 vs.
-0.99), but often fails to adequately reflect key fidelity markers such as distress monitoring.
-This work highlights gaps in current evaluation frameworks and advocate for fidelity-aware metrics that go beyond surface fluency to uncover clinically significant failures.
-Our findings clarify where synthetic data can effectively complement real-world datasets—and where critical limitations remain.
+Synthetic data adoption in healthcare is driven by privacy concerns, data access limitations, and high annotation costs.
+We explore synthetic Prolonged Exposure (PE) therapy conversations for PTSD as a scalable alternative for training clinical models.
+We systematically compare real and synthetic dialogues using linguistic, structural, and protocol-specific metrics like turn-taking and treatment fidelity.
+We introduce and evaluate PE-specific metrics, offering a novel framework for assessing clinical fidelity beyond surface fluency.
+Our findings show that while synthetic data successfully mitigates data scarcity and protects privacy, capturing the most subtle therapeutic dynamics remains a complex challenge.
+Synthetic dialogues successfully replicate key linguistic features of real conversations, for instance, achieving a similar Readability Score (89.2 vs.
+88.1), while showing differences in some key fidelity markers like distress monitoring.
+This comparison highlights the need for fidelity-aware metrics that go beyond surface fluency to identify clinically significant nuances.
+Our model-agnostic framework is a critical tool for developers and clinicians to benchmark generative model fidelity before deployment in sensitive applications.
+Our findings help clarify where synthetic data can effectively complement real-world datasets, while also identifying areas for future refinement.
 How Real Are Synthetic Therapy Conversations?
@@
 Wiese4  Saeed Abdullah1 1College of Information Sciences and Technology, Penn State University, USA 2Department of Psychiatry and Behavioral Sciences, Emory University, USA 3School of Interactive Computing, Georgia Tech, USA 4School of Psychology, Georgia Tech, USA {bnsuhas,saeed}@psu.edu, andrew.m.sherrill@emory.edu
+Figure 1: Selected annotated examples from synthetic PE therapy sessions.
+Each dialogue segment is evaluated for fidelity adherence based on PE protocol guidelines.
+Despite structural fluency, subtle violations like role drift (Ex.
+1) and premature reflection (Ex.
+4) highlight limitations in fidelity that escape automated scoring.
 1 Introduction
 Training machine learning models in sensitive domains like healthcare remains a challenge (Giuffrè and Shung, 2023; Kokosi and Harron, 2022).
-Access to real clinical conversations—crucial for modeling tasks like mental health diagnostics and therapeutic dialogue understanding—is severely limited by high annotation costs, patient privacy concerns (Kokosi and Harron, 2022; BN and Abdullah, 2022; BN et al., 2023), and ethical constraints on data sharing.
+Access to real clinical conversations, which are crucial for modeling tasks like mental health diagnostics and therapeutic dialogue understanding, is severely limited by high annotation costs, patient privacy concerns (Kokosi and Harron, 2022; BN and Abdullah, 2022; BN et al., 2023), and ethical constraints on data sharing.
 Synthetic data has emerged as a promising alternative, offering scalability and privacy preservation while reducing dependence on real-world annotations (Aher et al., 2023).
 In trauma-focused mental health care, particularly Prolonged Exposure (PE) therapy for PTSD, large language models (LLMs) can generate synthetic therapy dialogues at scale.
-However, questions remain about whether these dialogues capture more than surface-level fluency—specifically, the subtle dynamics of therapeutic fidelity such as emotional pacing, avoidance management, and protocol adherence (Shen et al., 2024).
-To address this, we develop and validate methods to measure clinically-relevant fidelity in generated dialogues—moving beyond standard metrics like coherence or perplexity.
-Prior work shows that while synthetic PE sessions can convincingly mimic real sessions in tone and structure (BN et al., 2025), they may still commit fidelity lapses, such as premature reflection or reinforcement of avoidance, which often go unnoticed by both automatic metrics and non-clinical annotators (Chiu et al., 2024; Zhang et al., 2024; Lee et al., 2024b).
+However, questions remain about whether these dialogues capture more than surface-level fluency, specifically the subtle dynamics of therapeutic fidelity such as emotional pacing, avoidance management, and protocol adherence (Shen et al., 2024).
+To address this, we develop and validate methods to measure clinically relevant fidelity in generated dialogues, moving beyond standard metrics like coherence or perplexity.
+Prior work shows that while synthetic PE sessions can convincingly mimic real sessions in tone and structure (BN et al., 2025b), they may still commit fidelity lapses, such as premature reflection or reinforcement of avoidance, which often go unnoticed by both automatic metrics and non-clinical annotators (Chiu et al., 2024; Zhang et al., 2024; Lee et al., 2024b).
 For instance, a therapist saying “That’s a really powerful insight” mid-exposure may appear empathic but violates PE protocol by derailing trauma processing (see Fig.
@@
 Without clinical expertise, both humans and automated metrics tend to overestimate fidelity.
-We introduce a fidelity-focused lens that evaluates dialogues on multiple dimensions: linguistic coherence, adherence to PE protocols, and the therapist’s navigation of key clinical interactions (e.g., managing avoidance, SUDS monitoring).
+We introduce a fidelity-focused lens that evaluates dialogues on multiple dimensions: linguistic coherence, adherence to PE protocols, and the therapist’s navigation of key clinical interactions (e.g., managing avoidance, SUDS (Subjective Units of Distress) monitoring).
 This framework integrates automated scoring and expert clinical assessment for a more rigorous evaluation of synthetic dialogue quality.
-We present the first large-scale, multi-dimensional comparison of real and synthetic PE therapy sessions—analyzing their linguistic, structural, and clinical fidelity characteristics.
+We present the first large-scale, multi-dimensional comparison of real and synthetic PE therapy sessions, analyzing their linguistic, structural, and clinical fidelity characteristics.
 Our findings demonstrate the current strengths and limitations of synthetic dialogues and inform future improvements in generation and evaluation methods.
-Our study contributes to bridging the real–synthetic data gap and provides a roadmap for advancing synthetic dialogue in sensitive healthcare domains.
+Our study contributes to bridging the real-synthetic data gap and provides a roadmap for advancing synthetic dialogue in sensitive healthcare domains.
 2 Motivation
@@
 Other studies focus on counselor style (Xie et al., 2024) and data augmentation (Kim et al., 2024).
-While Thousand Voices of Trauma introduces structural variation (BN et al., 2025), most evaluations compare to general counseling or CBT (Lee et al., 2024a; Shen et al., 2024) and rarely assess trauma-specific markers or PE’s structural fidelity, such as avoidance handling and imaginal exposure sequences.
+While The Pursuit of Empathy highlights the importance of empathy-aware evaluation (BN et al., 2025a), Thousand Voices of Trauma introduces structural variation (BN et al., 2025b), but most evaluations compare to general counseling or CBT (Lee et al., 2024a; Shen et al., 2024) and rarely assess trauma-specific markers or PE’s structural fidelity, such as avoidance handling and imaginal exposure sequences.
 Our work directly addresses this by evaluating whether synthetic PE sessions capture clinical realism per PE protocols (e.g., SUDS, avoidance redirection), highlighting procedural errors missed by standard metrics.
@@
 Psychological measures (CTRS, PANAS) in COUNSELINGEVAL assess empathy and coherence (Lee et al., 2024a), and interactional features like PQA are explored (Shen et al., 2024).
-However, most metrics are not adapted for trauma therapy, and methods like Dynamic Time Warping (DTW)—potentially useful for structured PE flows—remain underused.
+However, most metrics are not adapted for trauma therapy, and methods like Dynamic Time Warping (DTW), which are potentially useful for structured PE flows, remain underused.
 While some qualitative reviews consider emotional tone and goal alignment (Zhang et al., 2024; Xie et al., 2024), and structured questioning is explored (Ren et al., 2024), robust frameworks for quantitative evaluation of PE-specific structure are lacking.
@@
 Despite progress in synthetic dialogue generation (Qiu and Lan, 2024; Zhang et al., 2024; Xie et al., 2024), evaluation metrics (Lee et al., 2024a; Zhang et al., 2024; Xie et al., 2024), and human validation, a gap remains in fidelity assessment for trauma-focused therapies like PE (Chiu et al., 2024; Qiu and Lan, 2024; Lee et al., 2024a; Shen et al., 2024; Zhang et al., 2024; Xie et al., 2024; Hu et al., 2024).
-Current approaches lack metrics and focus tailored to PE’s unique conversational structure—including avoidance handling, trauma cue processing, and imaginal exposure flow—and do not address challenges such as patient emotional dysregulation (Qiu and Lan, 2024; Lee et al., 2024a; Shen et al., 2024; Zhang et al., 2024; Xie et al., 2024; Kim et al., 2024).
+Current approaches lack metrics and focus tailored to PE’s unique conversational structure, including avoidance handling, trauma cue processing, and imaginal exposure flow, and they do not address challenges such as patient emotional dysregulation (Qiu and Lan, 2024; Lee et al., 2024a; Shen et al., 2024; Zhang et al., 2024; Xie et al., 2024; Kim et al., 2024).
 Our work systematically evaluates synthetic data for alignment with PE structure and clinical validity, identifying areas for improving synthetic PE dialogue quality (see Table 1, Appendix A).
@@
 4 Methodology
-We analyzed 400 prolonged exposure therapy conversations—200 real-world and 200 synthetic.
-The synthetic data comes from the publicly available Thousand Voices of Trauma dataset (BN et al., 2025), generated using Claude Sonnet 3.5 (Anthropic, 2024) with PE-specific prompting frameworks adapted from clinical guidelines.
-This data includes diverse therapist-client interactions across trauma types, therapy phases, and demographics.
+We analyzed 400 prolonged exposure therapy conversations, consisting of 200 real-world and 200 synthetic.
+To demonstrate our evaluation framework, this case study uses 200 synthetic conversations from the publicly available Thousand Voices of Trauma dataset (BN et al., 2025b).
+The data was generated using Claude Sonnet 3.5 (Anthropic, 2024) with PE-specific prompting frameworks adapted from clinical guidelines.
+This dataset includes diverse therapist-client interactions across trauma types, therapy phases, and demographics.
 The real-world PE sessions were collected under IRB-approved protocols, with participant consent, and cannot be released due to privacy constraints.
-Each session (1–1.5 hours) was transcribed using Amazon HealthScribe (Services, 2023), manually verified, and reformatted by merging consecutive speaker turns for readability.
+Each session (1-1.5 hours) was transcribed using Amazon HealthScribe (Services, 2023), manually verified, and reformatted by merging consecutive speaker turns for readability.
 Each conversation followed a validated therapy fidelity checklist (See Table 1) to align with real-world standards.
@@
 These metrics provide insights into conversational dynamics, protocol adherence, and overall fidelity, ensuring a holistic evaluation of synthetic dialogue generation.
-To reduce subjectivity, we adopt a fidelity checklist (Table 1) and PE-specific metrics (Sec.
-5) grounded in existing clinical guidelines for evaluating therapeutic adherence.
+To reduce subjectivity, we adopt a fidelity checklist (Table 1) and PE-specific metrics grounded in existing clinical guidelines for evaluating therapeutic adherence.
 Our methodology consists of four key analyses: system-level metrics comparison, correlation analysis, statistical significance testing, and PE-specific metrics.
@@
     Predictability of text: Flow Entropy and Perplexity can help measure randomness and fluency, determining whether the synthetic text is overly structured or unnatural.
-    By comparing means and standard deviations, we evaluate whether synthetic conversations approximate real dialogue structures.
-Significant deviations signal potential areas where synthetic data falls short.
 4.1.2 Correlation Analysis
 While mean comparisons provide a general overview, correlation analysis (see Table 2) quantifies the consistency of relationships across metrics between real and synthetic data.
-This determines if real-data patterns persist in synthetic dialogues.
-For instance:
-  1. 1.
-    Turn-taking consistency: Strong correlations in Normalized Therapist/Client Turns and Turn Ratios indicate that synthetic dialogues replicate real-world conversational structures.
-  2. 2.
-    Linguistic coherence: A high correlation in Utterance Length and Flow Entropy suggests that synthetic responses maintain natural variation in sentence length and randomness.
-Lower correlation values suggest that synthetic data may lack fidelity in preserving conversational nuances, particularly in aspects such as vocabulary richness and readability.
 4.1.3 Statistical Significance Testing
 To confirm whether differences between real and synthetic data are statistically meaningful, we conduct Mann-Whitney U tests (Table 3).
-This assesses if variations are random or due to inherent inconsistencies.
-Key takeaways include:
-  1. 1.
-    Significant differences in linguistic complexity: Metrics such as Utterance Length, Turn Duration, and Vocabulary Richness show statistically significant differences, indicating that synthetic dialogues do not yet fully capture the expressive range of real conversations.
-  2. 2.
-    Minimal differences in turn-taking ratios: Therapist-Client Turn Ratio and Speaker Switches show weaker statistical significance, suggesting that structural patterns are relatively well-preserved.
-  3. 3.
-    Higher variability in generated text: The difference in Perplexity and Flow Entropy highlights challenges in maintaining linguistic variability without excessive repetition or over-simplification.
-Figure 1: Selected annotated examples from synthetic PE therapy sessions.
-Each dialogue segment is evaluated for fidelity adherence based on PE protocol guidelines.
-Despite structural fluency, subtle violations like role drift (Ex.
-1) and premature reflection (Ex.
-4) highlight limitations in fidelity that escape automated scoring.
+To identify which properties most strongly differentiate real and synthetic sessions, we next examine feature importance from a simple predictive model (see Table 4).
 4.1.4 PE-Specific Metrics
@@
 Manual fidelity annotations were then reviewed by a licensed clinical psychotherapist for adherence to PE protocol.
-While inter-rater agreement was not computed in this exploratory phase, these annotations served as a qualitative tool to identify and illustrate typical fidelity lapses—especially those potentially overlooked by automated metrics, rather than as quantitative ground truth.
+While inter-rater agreement was not computed in this exploratory phase, these annotations served as a qualitative tool to identify and illustrate typical fidelity lapses, especially those potentially overlooked by automated metrics, rather than as quantitative ground truth.
 Representative examples of such lapses identified through this process are illustrated in Figure 1 and discussed in Section 6.2.
@@
   2. 2.
-    Concise and structured responses: Synthetic dialogues are shorter (Average Utterance Length: 22.90 ± 1.74 vs.
-68.72 ± 26.61) and more consistent (Utterance Length Std Dev: 18.54 ± 2.35 vs.
-135.85 ± 66.25) due to LLM output constraints.
-Techniques like Chain-of-Thought prompting helps improve coherence, though larger output contexts are needed.
+    Concise and structured responses: Synthetic dialogues are shorter (Average Utterance Length: 22.90 $\pm$ 1.74 vs.
+68.72 $\pm$ 26.61) and more consistent (Utterance Length Std Dev: 18.54 $\pm$ 2.35 vs.
+135.85 $\pm$ 66.25) due to LLM output constraints.
+Techniques like Chain-of-Thought prompting help improve coherence, though larger output contexts are needed.
   3. 3.
@@
   5. 5.
-    Correlation analysis: High correlations in turn-taking metrics (0.65–0.78 synthetic vs.
-0.85–0.92 real) confirm accurate conversational dynamics.
+    Correlation analysis: High correlations in turn-taking metrics (0.65-0.78 synthetic vs.
+0.85-0.92 real) confirm accurate conversational dynamics.
 Lower correlations in linguistic complexity metrics (e.g., Readability Score, Vocabulary Richness, Perplexity) reflect structural differences but do not hinder the coherence/naturalness of the conversation.
@@
 Norm. Spkr. Switches $0.99\pm 0.0$ $0.98\pm 0.0$ $0.85$ $0.72$
-Norm. Total Turns $1.00\pm 0.0$ $1.00\pm 0.0$ $0.78$ $0.65$
+Norm. Total Turns $\pm 0.0 1.00±0.01$$\pm 0.0$ $0.78$ $0.65$
 Avg. Utt. Len. $68.7\pm 26.6$ $22.9\pm 1.7$ $0.91$ $0.76$
@@
 Norm. C Words $46.8\pm 22.5$ $18.0\pm 1.7$ $0.90$ $0.75$
-Turn Ratio (T/C) $0.01\pm 0.0$ $0.01\pm 0.0$ $0.80$ $0.66$
-Word Ratio (T/C) $0.01\pm 0.0$ $0.00\pm 0$ 0.0 $0.82$ $0.68$
+Turn Ratio (T/C) $01\pm 0.0 0.0$$01\pm 0.0$ $0.80$ $0.66$
+Word Ratio (T/C) $1\pm 0.0 0.00±0$$0\pm 0.0$ $0.82$ $0.68$
 Vocab. Richness $0.13\pm 0.0$ $0.18\pm 0.0$ $0.77$ $0.63$
@@
 This suggests that while individual responses may be more concise, the overall flow and engagement in the conversation are well-preserved.
+For example, while a real client might repeatedly use an informal phrase like “freaked out,” the LLM generates diverse but semantically similar alternatives such as “felt anxious,” “was overcome with fear,” or “grew very distressed,” thereby increasing lexical richness within a tight, clinically-appropriate boundary.
 These findings reinforce that the synthetic model captures key conversational characteristics while ensuring structured, coherent responses.
@@
 Norm. C Words $2.47\times 10^{3}$ $p<0.001$
-T-C Turn Ratio $2.13\times 10^{3}$ $1.22\times 10^{-9}$
+T-C Turn Ratio $3\times 10^{3} 1.22×1$$2\times 10^{-9}$
 T-C Word Ratio $2.37\times 10^{3}$ $p<0.001$
@@
 5.3 Feature Importance
-Feature importance analysis identifies the metrics most influencing synthetic data generation, confirming trends from previous evaluations.
+Feature importance analysis (see Table 4) identifies the metrics most influencing synthetic data generation, confirming trends from previous evaluations.
 While not the central focus, these results support the model’s alignment with conversational structure and linguistic patterns.
@@
   3. 3.
-    Lexical richness is moderately important: Vocabulary Richness plays a secondary role, reflecting that synthetic dialogues maintain varied but structured language.
-While real conversations are more flexible, synthetic responses balance vocabulary diversity with clarity.
+    Lexical richness is moderately important: Synthetic dialogues use varied yet structured language, balancing vocabulary diversity with clarity, while real conversations are more flexible.
 5.4 Prolonged Exposure Specific Metrics
 Table 5 presents the statistical test results comparing synthetic Prolonged Exposure data with real therapy session data.
-The results indicate that the synthetic data effectively captures key therapeutic constructs, with several core PE metrics such as Trauma Narrative Coherence, Emotional Engagement, Avoidance Handling, Exposure Guidance, Cognitive Restructuring, Avoidance Reduction, and Emotion Intensity showing strong alignment with real data (p < 0.001) showing strong alignment with real data.
-These results suggest that the synthetic data successfully replicate the structure and engagement patterns seen in real therapy sessions.
-However, certain metrics do not reach statistical significance, indicating areas where synthetic data generation may need refinement.
-Emotional Habituation ($p=0.102$) and SUDS Progression ($p=0.073$) track distress reduction over repeated exposures, and their lack of significance suggests the synthetic data may not fully model this progressive decrease.
-Similarly, while Trauma Narrative Coherence is well-captured, Narrative Development ($p=0.251$) appears to be less robust, possibly due to limited context retention or rigid structure in generated responses.
+The synthetic data shows promise, with no statistically significant difference from real sessions on key temporal metrics like Emotional Habituation ($p=0.102$) and SUDS Progression ($p=0.073$).
+Similarly, Narrative Development ($p=0.251$) appears less robust but also shows no significant statistical difference, possibly due to limited context retention.
+This suggests the model is beginning to capture the therapeutic arc of distress reduction.
+Conversely, foundational constructs such as Trauma Narrative Coherence, Emotional Engagement, and Avoidance Handling show statistically significant differences (p < 0.001) when compared to real dialogues.
+These findings provide a clear and actionable roadmap for future work, guiding efforts to calibrate the model’s consistency with the variability of real-world therapeutic interactions.
 Despite these discrepancies, the overall performance of synthetic data suggests it is a viable alternative for privacy-sensitive applications.
@@
 First, synthetic PE sessions can serve as valuable training tools for novice therapists learning to identify protocol deviations, as they offer controlled examples of both adherent and non-adherent interactions without privacy concerns.
+Crucially, this fidelity-focused evaluation serves as an essential ethical prerequisite.
+Before any AI tool for therapy can be considered for patient-facing studies, a framework like ours must be used to verify its adherence to clinical protocols and prevent the risk of clinically significant failures.
 Second, the consistent replication of structural elements (e.g., turn-taking, session flow) indicates that synthetic data can effectively supplement clinical training materials, particularly in settings with limited access to specialized trauma training.
 However, our qualitative analysis reveals limitations that automated metrics may overlook.
-As shown in Figure 1, even linguistically fluent synthetic dialogues can contain subtle but clinically significant fidelity violations—including role drift, premature processing, and improper SUDS implementation—that typically escape quantitative detection.
+As shown in Figure 1, even linguistically fluent synthetic dialogues can contain subtle but clinically significant fidelity violations, such as role drift, premature processing, and improper SUDS implementation, that typically escape quantitative detection.
 These lapses would likely go unnoticed by non-specialist reviewers, highlighting the necessity of expert-guided evaluation for AI tools in clinical contexts.
@@
 These specific instances demonstrate common patterns of deviation from PE protocol:
-  • Ex. 1: Role Drift — The therapist assumes the client’s narrative responsibility, undermining the therapeutic goal of client-driven processing.
-  • Ex. 2: Failure to Maintain Focus on Exposure Targets — Avoidance behavior is permitted, a lapse not distinguishable by standard metrics.
-  • Ex. 3–4: Premature Reflection and Intervention — The therapist interjects before full emotional processing, disrupting the habituation process crucial for PE efficacy.
-  • Ex. 5: Improper SUDS Implementation — The dialogue maintains structural flow, but the clinical assessment validity is compromised.
+  • Fidelity Lapse (Ex. 1): Role Drift - The therapist reflects on the client’s observation instead of redirecting them to the trauma narrative, which breaks the PE protocol mid-exposure.
+  • Fidelity Issue (Ex. 2): Generic Affirmation - The therapist uses an overused phrase that lacks the specific grounding necessary for an anxious, compulsive client.
+  • Protocol Adherence (Ex. 3): No Issue - The client’s dialogue shows they are appropriately immersed in the trauma memory, which is the goal for imaginal exposure.
+  • Fidelity Lapse (Ex. 4): Reflection During Exposure - The client prematurely shifts into cognitive analysis during the exposure, a process that belongs in post-session processing.
+  • Protocol Adherence (Ex. 5): Trauma Anchoring - The therapist correctly redirects the client back to the trauma scene after a drift, which successfully maintains the integrity of the PE protocol.
 These detailed examples underscore why evaluation frameworks must integrate clinical expertise alongside computational metrics to accurately assess therapeutic fidelity in synthetic PE sessions.
@@
 We contribute a fidelity-aware evaluation framework tailored to Prolonged Exposure (PE) therapy, along with metrics highlighting that current generative models, despite their fluency, lack the nuance needed for high-stakes therapeutic contexts.
+The metrics and methods in this framework are model-agnostic and can be readily applied by the research community to benchmark and improve dialogues generated from any large language model.
 Our findings underscore the risk of overestimating quality through non-expert or automated evaluations.
@@
 For details on the ethical safeguards, simulation design, and usage guidelines related to the synthetic dataset, we refer readers to BN et al.
-(2025).
+(2025b).
+Acknowledgement
+This work was supported by the National Science Foundation (NSF) under Grant No.
+2326144.
+Any opinions, findings, and conclusions or recommendations expressed are those of the author(s) and do not necessarily reflect the views of the NSF.
+10 Data Availability
+The synthetic dataset (Thousand Voices of Trauma) used in this paper is publicly available at https://huggingface.co/datasets/yenopoya/thousand-voices-trauma under the CC BY-NC 4.0 license.
+For collaborating on the real-world dataset, please contact PI Dr.
+Sherrill.
 Limitations
-While our study demonstrates the promise of synthetic dialogues in approximating real PE therapy interactions, several limitations remain.
-Although we measure fidelity using structural, statistical, and protocol-based metrics, our evaluation does not assess therapeutic effectiveness or downstream clinical outcomes.
-Synthetic data that aligns structurally with real dialogues may still fall short in supporting meaningful therapeutic engagement or behavior change, especially in high-stakes or emotionally complex scenarios.
-Moreover, we do not evaluate inter-rater agreement on fidelity violations, which is critical for establishing the robustness of manual annotations—this is planned for future work.
-Acknowledgement
-This work is supported by the National Science Foundation (NSF) under Grant No.
-(Details omitted for double-blind review.) Any opinions, findings, & conclusions or recommendations expressed in this work are those of the author(s) and do not necessarily reflect the views of the NSF.
+Although our study demonstrates the promise of synthetic dialogues in approximating real PE therapy interactions, several limitations remain: (1) our analysis is presented as a deep case study focused on data generated by a single, state-of-the-art LLM; (2) future work should apply our framework to compare outputs from a wider range of models to assess the generalizability of these specific findings; (3) although we measure fidelity using structural, statistical, and protocol-based metrics, our evaluation does not assess therapeutic effectiveness or downstream clinical outcomes; (4) synthetic data that aligns structurally with real dialogues may still fall short in supporting meaningful therapeutic engagement or behavior change, especially in high-stakes or emotionally complex scenarios; (5) we do not evaluate inter-rater agreement on fidelity violations, which is critical for establishing the robustness of manual annotations, and this is planned for future work.
 Appendix A Metric Definitions